### v1

# Uncovering Bias Paths with LLM-guided Causal Discovery: An Active Learning and Dynamic Scoring Approach

Abstract
Causal discovery (CD) plays a pivotal role in understanding the mechanisms underlying complex systems. While recent algorithms can detect spurious associations and latent confounding, many struggle to recover fairness-relevant pathways in realistic, noisy settings. Large Language Models (LLMs), with their access to broad semantic knowledge, offer a promising complement to statistical CD approaches, particularly in domains where metadata provides meaningful relational cues. Ensuring fairness in machine learning requires understanding how sensitive attributes causally influence outcomes, yet CD methods often introduce spurious or biased pathways. We propose a hybrid LLM-based framework for CD that extends a breadth-first search (BFS) strategy with active learning and dynamic scoring. Variable pairs are prioritized for LLM-based querying using a composite score based on mutual information, partial correlation, and LLM confidence, improving discovery efficiency and robustness.
To evaluate fairness sensitivity, we construct a semi-synthetic benchmark from the UCI Adult dataset, embedding a domain-informed causal graph with injected noise, label corruption, and latent confounding. We assess how well CD methods recover both global structure and fairness-critical paths (e.g., sex $\rightarrow$ education $\rightarrow$ income).
Our results show that LLM-guided methods, including the proposed method, demonstrate competitive or superior performance in recovering such pathways under noisy conditions. We highlight when dynamic scoring and active querying are most beneficial and discuss implications for fairness auditing in real-world datasets.
1 Introduction
Causal discovery (CD) identifies causal relationships among variables and is widely used in domains such as epidemiology, genetics, and economics (Kampani et al. 2024). Recent advances in CD algorithms have opened new avenues for addressing bias and fairness in machine learning (ML). Bias in ML often arises from hidden confounders or unequal treatment of sensitive groups (Barocas, Hardt, and Narayanan 2023). For instance, a hiring algorithm may inherit biased patterns from historical data that unfairly disadvantage certain demographic groups (Binns 2018). CD helps identify such pathways, distinguishing genuine effects from biases introduced by confounders and ensuring decisions are based on causal factors rather than proxies (Makhlouf, Zhioua, and Palamidessi 2020). Unlike traditional methods relying on statistical correlations or predefined fairness metrics, CD disentangles direct effects, indirect effects, and spurious correlations (Miguel-Hurtado et al. 2016).
This has prompted growing interest in using CD to support fairness-aware ML. Recent studies have proposed using counterfactual reasoning, structural causal models (SCMs), and fairness-specific constraints to detect and mitigate discriminatory pathways (Binkytė et al. 2023; Zuo et al. 2022; Wang and Luo 2023). However, classical CD methods often rely on strong assumptions, perform poorly under noise or confounding, and struggle to align with domain knowledge not explicitly encoded in data (Takayama et al. 2024).
Large Language Models (LLMs) have emerged as promising tools for CD, using their vast semantic knowledge to infer causal directions between variables such as “sex” and “income” from textual metadata (Le, Xia, and Chen 2024; Vashishtha et al. 2023; Khatibi et al. 2024; Kampani et al. 2024; Long, Schuster, and Piché 2023; Kıcıman et al. 2023). Jiralerspong et al. (Jiralerspong et al. 2024) proposed a breadth-first search (BFS) strategy for efficient graph construction via LLM queries, showing improved performance over random or exhaustive querying. However, a growing concern is that CD methods, especially when used in fairness audits, may introduce or amplify bias by inferring spurious links involving sensitive features. In fairness-sensitive contexts, incorrect causal attributions can lead to unjustified conclusions about discrimination or fairness violations. Thus, ensuring both the accuracy and interpretability of discovered pathways is essential for responsible ML deployment.
Building on prior work, we introduce an enhanced BFS framework that integrates Active Learning (AL) and Dynamic Scoring. Instead of querying all pairs uniformly, our method ranks pairs using mutual information (MI), partial correlation (PCorr), and LLM-derived confidence, while discouraging redundant queries and focusing on informative regions via a history-based weight. This approach treats the LLM not as a passive responder but as an adaptive component within an iterative learning loop.
A key challenge in evaluating CD methods in fairness contexts is the absence of real-world datasets with known ground-truth causal structures involving sensitive attributes (Loftus et al. 2018). Most, like the UCI Adult dataset, offer rich features but lack verified causal graphs. This limitation makes it difficult to validate whether inferred pathways truly reflect bias mechanisms or are artifacts of the discovery process. To address this, we construct a semi-synthetic benchmark that retains real-world semantics while embedding a known fairness-relevant graph for controlled evaluation.
We analyze the performance of our proposed method and other baseline CD methods on the constructed semi-synthetic benchmark based on the UCI Adult dataset that incorporates fairness-relevant structure, noise, and latent confounding. We evaluate these methods across their ability to recover key pathways (e.g., sex $\rightarrow$ education $\rightarrow$ income) and to support fairness audits.
We also conduct a hyperparameter sensitivity analysis to examine how LLM-specific factors such as temperature, query prioritization weights, and stopping criteria affect CD performance. These findings offer practical insights into how LLM behavior and prompting influence graph reconstruction accuracy and fairness-path recovery.
Our contributions are as follows:
  • We propose a hybrid LLM-based CD method that integrates AL and dynamic scoring to prioritize informative queries during BFS-based graph recovery.
  • We introduce a semi-synthetic benchmark grounded in fairness literature, with latent confounding, label noise, and sensitive pathways, to evaluate CD methods under realistic conditions.
  • We provide a comparative analysis of pathway recovery across methods and examine how different discovery strategies affect fairness-related inferences.
  • We analyze the impact of LLM-specific hyperparameters on discovery quality.
2 Related Work
Causal discovery has been explored through statistical, optimization-based, and more recently, LLM-guided approaches (Long et al. 2025).
Classical methods like the Peter-Clark (PC) algorithm (Spirtes and Glymour 1991) and Greedy Equivalence Search (GES) (Meek 1997) use conditional independence tests or score-based search but rely on strong assumptions (e.g., causal sufficiency, faithfulness) and degrade with latent confounding or limited data. Optimization-based approaches, such as NOTEARS (Zheng et al. 2018) and DAGMA (Bello, Aragam, and Ravikumar 2022), frame CD as a constrained optimization problem. While accurate, they are computationally expensive and sensitive to hyperparameters, limiting their use in fairness-critical or iterative settings.
LLM-based CD methods offer an alternative by leveraging variable descriptions to infer structure, offering a way to encode human knowledge and priors. Kıcıman et al. (Kıcıman et al. 2023) introduced pairwise LLM queries using only metadata, though the quadratic query complexity limits scalability. Others combine LLM-derived causal orders with algorithms like PC (Vashishtha et al. 2023), or embed LLMs in multi-agent or prompt-refinement loops (Khatibi et al. 2024; Le, Xia, and Chen 2024). Kampani et al. (Kampani et al. 2024) take a hybrid approach, using LLMs to generate priors, refining them with methods like NOTEARS.
Takayama et al. (Takayama et al. 2024) propose statistical causal prompting, where LLMs enhance statistical methods by providing priors or plausible causal orders. While robust, this one-time augmentation lacks iterative refinement. Similarly, Jiralerspong et al. (Jiralerspong et al. 2024) propose a BFS framework to reduce LLM query complexity during structure discovery. Their approach treats LLM responses as inputs for efficient edge discovery, dramatically lowering runtime without sacrificing metadata grounding. However, BFS applies fixed selection heuristics and does not dynamically balance statistical evidence with LLM insights, making it susceptible to early decision errors.
Our method builds on this BFS foundation, introducing AL and a dynamic scoring strategy that prioritizes variable pairs based on a mixture of MI, PCorr, and LLM confidence scores. Compared to prior methods that treat all variable pairs uniformly or rely solely on LLMs or statistics, our approach dynamically balances complementary signals and avoids over-reliance on any single information source.
Some of the above-mentioned approaches differ fundamentally from our framework in both scope and evaluation goals. Methods such as those proposed by Vashishtha et al. (Vashishtha et al. 2023), Khatibi et al. (Khatibi et al. 2024), Le et al. (Le, Xia, and Chen 2024), and Kampani et al. (Kampani et al. 2024) use LLMs to generate global causal orders or priors that are later refined using constraint-based or optimization-based algorithms like PC or NOTEARS. In contrast, our method treats the LLM as an active querying agent embedded in an iterative loop, where causal judgments are elicited at the pairwise level and guided dynamically by a scoring function informed by both statistical and historical signals. Unlike prompting-based or prior-initialized approaches, our framework does not rely on a static, up-front extraction of priors, but instead continuously adapts query selection through AL.
We further emphasize fairness-aware causal recovery, targeting sensitive paths and mediators (e.g., sex $\rightarrow$ education $\rightarrow$ income). This responds to calls for causal approaches to fairness (Nabi and Shpitser 2018; Kilbertus et al. 2017; Zhang and Bareinboim 2018; Chiappa 2019), highlighting the need for explainable, intervenable, and robust fairness assessments under noisy conditions. In this sense, our method contributes to a growing understanding of how LLM-based querying, when combined with statistical reasoning, can affect fairness-relevant causal inference under noisy and complex conditions.
3 Proposed LLM-Based Method
3.1 LLM-Guided Breadth-First Search for Causal Discovery
The foundation of the proposed method is the BFS-based CD framework introduced by (Jiralerspong et al. 2024). This approach iteratively queries an LLM to infer causal relationships among variables in a data set and reduces the number of queries required since visiting each node once is sufficient to determine all edges originating from that node. Using this prompting approach, each query is approached as an opportunity for node expansion within a BFS algorithm, progressively building the causal graph as it is traversed using BFS. To effectively implement the BFS, it is essential to determine the sequence in which nodes are explored. For this purpose, the Directed Acyclic Graph (DAG) nature of causal graphs is leveraged, by using their topological ordering to guide the BFS traversal. The BFS process follows three main steps:
  1. 1.
    Initialization: The algorithm begins by querying the LLM to identify independent variables that are not caused by any other variables which form the starting points for graph exploration.
  2. 2.
    Expansion: For each variable on the current frontier (independent variables or those already explored), the LLM is queried about which other variables it causally influences, using prompts that incorporate discovered structure so far.
  3. 3.
    Graph Construction: Inferred edges are added to a directed graph if they do not introduce cycles. Each new relationship is checked to ensure that the variables in the frontier are connected to previously unexplored nodes. The process continues until all nodes are explored. The final output is a predicted causal graph in the form of an adjacency matrix.
While more efficient than exhaustive pairwise querying, this approach still explores variable pairs uniformly. As a result, many LLM queries target uninformative pairs, especially in sparse or noisy graphs, motivating the need for adaptive query selection.
3.2 Active Learning and Dynamic Scoring for Efficient Querying
To improve the informativeness of queries and reduce redundancy, we introduce a scoring-based AL framework. This component augments the BFS approach with a dynamic prioritization mechanism for selecting the most valuable variable pairs to query. Figure 1 illustrates the full framework.
Figure 1: Overview of the proposed LLM-guided BFS framework with dynamic scoring and AL.
Dynamic Scoring Mechanism
For each unqueried pair $(x,y)$, a dynamic informativeness score $S(x,y)$ is computed as:
$\begin{split}S(x,y)=\ &w_{\text{stat}}\cdot\text{StatScore}(x,y)+\\ &w_{\text{conf}}\cdot\text{LLMConf}(x,y)+w_{\text{hist}}\cdot\text{HistScore}(x,y)\end{split}$ (1)
Each term incorporates a distinct signal:
  • Statistical Score: Combines mutual information (MI) and partial correlation (PCorr), measuring both linear and non-linear dependencies:
    $\text{StatScore}(x,y)=\frac{\text{MI}(x,y)+\text{PCorr}(x,y)}{2}$ (2)
  • LLM Confidence Score: Reflects the certainty of the model’s previous response for a given pair, where higher token-level confidence leads to higher scores:
    $\text{LLMConf}(x,y)=\frac{1}{1+e^{-\text{confidence}}}$ (3)
  • Query History Score: Penalizes repeated queries to encourage broader exploration:
    $\text{HistScore}(x,y)=\frac{1.5}{1+\text{query\_count}(x,y)}$ (4)
The weights $w_{\text{stat}},w_{\text{conf}},w_{\text{hist}}$ are treated as tunable hyperparameters. These weights are optimized via Bayesian optimization (see Section 7), allowing the framework to adapt scoring behavior to different graph structures or domain characteristics.
Query Selection via Active Learning
We incorporate dynamic scoring into an AL loop that determines which pairs to query. The query process iteratively selects:
$(x^{*},y^{*})=\arg\max_{(x,y)\in\text{UnqueriedPairs}}S(x,y)$ (5)
An LLM is then prompted to assess whether $x^{*}$ causally affects $y^{*}$ using the natural language format described in listing 1.
Listing 1 Causal query prompt format
Does X cause Y? Provide the answer in the format: <Answer>Yes</Answer> or <Answer>No</Answer>.
If the LLM predicts a positive causal link and the resulting graph remains acyclic, the edge $x^{*}\rightarrow y^{*}$ is added; otherwise, it is discarded. To estimate the LLM uncertainty, we extract token-level log-probabilities and compute the average probability of the model’s response (“Yes” or “No”). This score, scaled to [0,1], informs the dynamic scoring function. If log-probs are unavailable, a default confidence of 0.5 is used. The process terminates after a query limit or when scores fall below a threshold.
Cycle Detection and Output Construction
To ensure the output is a valid DAG, each edge addition is immediately followed by a cycle check. If the graph becomes cyclic, the addition is reversed.
The final causal graph is returned both as a directed edge list and an adjacency matrix:
$A(i,j)=\begin{cases}1&\text{if }X_{i}\rightarrow X_{j}\text{ is predicted},\\ 0&\text{otherwise}.\end{cases}$ (6)
The entire procedure is conducted in a multi-turn chat session, allowing the LLM to maintain contextual awareness of previously inferred relationships and rationales.
4 Fairness Evaluation through Causal Pathway Analysis
To evaluate the fairness utility of learned causal graphs, we perform pathway analysis to quantify how sensitive variables influence target outcomes. Unlike classical fairness metrics based on statistical associations (e.g., demographic parity), pathway analysis distinguishes between direct, mediated, and spurious effects, offering a principled causal perspective to fairness assessment (Pearl 2009, 2022).
We analyze the structure and influence of paths from sensitive attributes $S$ (e.g., race, sex) to outcome variables $Y$ (e.g., income) in the learned graphs. This approach distinguishes between direct, mediated, and spurious influences, going beyond traditional statistical fairness metrics that lack causal grounding (Pearl 2022).
Path Count Comparison.
For each discovered graph, we enumerate and classify all directed paths originating from sensitive variables:
  • Direct Paths ($P_{\text{direct, fairness}}$): Number of direct causal edges ($S\rightarrow Y$) that connect the defined sensitive variables to the outcomes.
  • Indirect Paths ($P_{\text{indirect, fairness}}$): Number of causal paths ($S\xrightarrow{}\text{intermediate nodes}\xrightarrow{}Y$) in which sensitive variables influence the outcomes indirectly via intermediate variables.
  • Spurious Paths ($P_{\text{spurious, fairness}}$): Number of paths that involve sensitive variables but do not influence the outcomes ($S\xrightarrow{}\text{non-$Y$ nodes}$).
To assess structural alignment with fairness-relevant pathways, we report the number of direct and indirect paths for each graph and compare these across CD methods and to the known ground-truth graph. A higher overlap with true fairness-relevant paths indicates better recovery of bias mechanisms.
Quantitative Effect Decomposition.
To assess the magnitude of fairness-related effects, we estimate:
  • Direct Effect (DE): The estimated causal effect of $S$ on $Y$ along direct paths.
  • Indirect Effect (IE): The effect of $S$ on $Y$ through mediating variables.
  • Total Effect (TE): The sum of direct and indirect effects:
    $TE=DE+IE$
These effects are computed using structural equations or interventional estimators (depending on the graph), and normalized to account for scale:
$C_{\text{bias}}=\frac{TE}{\text{Var}(Y)}$ (7)
where $C_{\text{bias}}$ represents the total fairness-relevant contribution of $S$ to variance in the outcome. This enables comparative evaluation across datasets and methods. It quantifies the total effect of sensitive attributes on the outcome (income), scaled by outcome variance. A higher $C_{\text{bias}}$ indicates a greater risk of propagating sensitive information through direct or mediated pathways, particularly important for fairness auditing.
5 Synthetic Graph Construction Based on UCI Adult Attributes
To evaluate fairness-aware CD under controlled yet realistic conditions, we construct a semi-synthetic dataset based on the UCI Adult dataset (Kohavi et al. 1996), commonly used in fairness research for its demographic diversity and income disparities. Our synthetic graph retains the original 15 variables and encodes plausible causal links informed by prior work (Kilbertus et al. 2017; Nabi and Shpitser 2018).
The DAG assumes that demographic features (e.g., race, sex) influence education and occupation; education affects income both directly and via mediators like occupation and capital gains; and age and marital status serve as confounders or mediators in employment-related paths. Capital-related features are modeled as downstream effects of occupational history, contributing to income.
To simulate structural unfairness, we inject direct edges from sex and race to income, consistent with causal fairness benchmarks (Nabi and Shpitser 2018; Kilbertus et al. 2017). These pathways are complemented by indirect paths (e.g., sex $\rightarrow$ education $\rightarrow$ occupation $\rightarrow$ income) to facilitate the estimation of direct, indirect, and total effects in fairness evaluation. The baseline graph and technical details on the data generation process can be found in Appendix A.1.
5.1 Structured Noise Augmentation
To test robustness under realistic conditions, we augment the clean synthetic data with structured noise and confounding, emulating sources of measurement error and label corruption common in real-world datasets.
Additive Noise for Continuous Variables.
Gaussian noise is added to all continuous variables post-generation to simulate measurement variability.
For each continuous variable X, noise is sampled as:
$X_{i}^{\text{noisy}}=X_{i}+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{X}^{2}),\quad\sigma_{X}=\alpha\cdot\mathrm{std}(X)$ (8)
where $X_{i}$ is the original value for sample $i$, $\epsilon_{i}$ is a random error drawn from a normal distribution with mean 0 and standard deviation $\sigma_{X}$, $\sigma_{X}=\alpha\cdot\mathrm{std}(X)$, where $\alpha\in[0,1]$ is a tunable noise scaling factor set to $\alpha=0.1$. This maintains proportionality between noise magnitude and feature variability, consistent with additive noise models used in ICA-LiNGAM and structural equation modeling (SEM) frameworks (Shimizu et al. 2006; Bollen 1989).
Symmetric Label Noise in Categorical Variables
To emulate mislabeling, we introduce symmetric label noise into categorical variables. For each categorical variable, a proportion $\gamma$ of values (5%) are randomly replaced with an alternative category, drawn uniformly from the remaining labels. For each categorical variable $Y$ with $K$ classes (e.g., ”Male”, ”Female”), uniform symmetric noise is introduced using the following equation.
Let $\gamma\in[0,1]$ be the label noise probability. Then for each sample $i$:
$P(\tilde{Y}_{i}=y^{\prime})=\begin{cases}1-\gamma&\text{if }y^{\prime}=Y_{i}\\ \frac{\gamma}{K-1}&\text{if }y^{\prime}\neq Y_{i}\end{cases}$ (9)
Where $Y_{i}$ is the original category label, $\tilde{Y}_{i}$ is the observed (possibly noisy) label, $K$ is the number of possible categories for $Y$, and $y^{\prime}\in\mathcal{Y}$ are all possible class values.
This means with probability $\gamma$, the label is flipped uniformly to one of the $K-1$ incorrect categories. And with probability $1-\gamma$, the label remains correct. This uniform corruption model known as symmetric label noise, is widely used in label-noise robust learning (Natarajan et al. 2013; Patrini et al. 2017).
Latent Confounding via Unobserved Variable $U$
To simulate unmeasured social influences (e.g., privilege or institutional access), we introduce a latent confounder $U\sim\mathcal{N}(0,1)$ that affects intermediate variables such as years of education, capital gain, and work hours, thereby indirectly influencing income. This reflects latent societal advantages that are not captured in observed features.
The latent confounder is injected via additive linear terms into structural equations following standard assumptions in structural causal models (SCMs). For the binary outcome income, we use a nonlinear threshold function that combines observable and latent contributions. This setup reflects bias amplification near decision boundaries and is consistent with prior fairness simulation frameworks. Full equations and implementation details are provided in Appendix A.2.
Controlled Replication Across Seeds
To evaluate generalization and fairness robustness, we generate four datasets using different random seeds, each with the same underlying graph but distinct noise realizations. Noise is added post-graph generation, preserving causal structure while introducing realistic variability. This mirrors standard practices in fairness benchmarking and robustness testing for CD.
6 Datasets
To evaluate the performance of the proposed LLM-based CD framework with AL and dynamic scoring, a series of experiments are conducted on benchmark networks. These networks are described in the following.
Child Causal Network.
The Child network is a medium-scale Bayesian network simulating causal relationships relevant to congenital heart disease in newborns (Spiegelhalter et al. 1993). It includes 20 nodes and 25 edges spanning medical, parental, and environmental factors. Variables include both categorical and continuous types, covering clinical features such as birth asphyxia, lung function, cardiac mix, chest X-ray results, and outcomes like mortality and hospital stay.
Neuropathic Pain Causal Network.
The Neuropathic Pain Diagnosis Simulator (Tu et al. 2019) models causal relationships underlying neuropathic pain using a graph of 221 nodes and 770 edges, based on real patient data. It includes symptom, pattern, and pathophysiological diagnoses from actual patient data.
As these benchmarks lacks sensitive variables, we do not assess fairness; evaluation focuses solely on structural accuracy.
7 Experiments
To evaluate the proposed LLM-based CD framework, we conduct a series of experiments on the introduced datasets. The primary goal is to assess the framework’s ability to improve causal graph accuracy and recover fairness-relevant pathways under varying configurations and noise conditions. This section outlines our experimental design, evaluation metrics, and baseline comparisons.
7.1 Baseline Methods
We compare our method against the following CD approaches:
Peter-Clark (PC) Algorithm.
A constraint-based method using conditional independence tests to iteratively prune edges from a fully connected graph, followed by orienting edges to form a DAG. It assumes causal sufficiency and the faithfulness assumption, relying heavily on accurate independence tests (Spirtes and Glymour 1991).
Greedy Equivalence Search (GES) Algorithm.
A score-based method that operates in two phases: a forward phase that incrementally adds edges to improve a score function (e.g., BIC), and a backward phase that removes edges to optimize the graph structure further. It assumes causal sufficiency and focuses on equivalence classes of DAGs to improve computational efficiency (Meek 1997).
NOTEARS Algorithm.
Formulates CD as a continuous optimization problem by introducing a differentiable acyclicity constraint. This allows it to learn DAGs using gradient-based optimization, balancing data fit and sparsity (Zheng et al. 2018).
DAGMA Algorithm.
Integrates neural networks and optimization techniques to model nonlinear causal relationships. It uses a masked autoencoder architecture to enforce sparsity and incorporates a differentiable acyclicity condition to ensure that the resulting graph is a valid DAG (Bello, Aragam, and Ravikumar 2022).
LLM-based Pairwise Method.
Utilizes LLMs to infer causal relationships through pairwise queries based on metadata and variable descriptions. While intuitive, it scales poorly with graph size (Kıcıman et al. 2023).
LLM-based BFS Method.
LLM-based BFS appraoch introduced by Jiralerspong et al. (Jiralerspong et al. 2024) which was discussed in the Sections 1 and 3.
All baselines are implemented using the open-source framework from (Jiralerspong et al. 2024) for consistency.
7.2 Experimental Design
We apply Bayesian optimization with a Gaussian Process (GP) surrogate model to explore the influence of key hyperparameters on performance. This enables tuning of our query scoring and LLM prompting strategy to enhance edge discovery and fairness-related signal recovery (Mockus 1994; Snoek, Larochelle, and Adams 2012). Optimization is performed using gp_minimize from scikit-optimize.
The hyperparameters optimized include:
  • Scoring Weights: Relative contributions of MI, PCorr, and query history in Equation 1.
  • Score threshold: Minimum dynamic score for querying a variable pair.
  • LLM Temperature: Controls the stochasticity of LLM responses.
  • Max Iterations: Maximum number of AL iterations per trial.
To evaluate robustness, we divide trials into chunks that sample different regions of the hyperparameter space. Each chunk records graph outputs and performance metrics, and the top-performing configuration from each run is used for final reporting. Full details on the hyperparameters used in each set of experiments are provided in Appendix B.
We evaluate all methods on four synthetic datasets generated from the Adult-based causal graph (Section 5) using different random seeds, which preserve the underlying structure while introducing independent data realizations. Performance metrics are averaged across these seeds. Fairness pathway analyses are conducted only on the Adult-based datasets, focusing on paths from sensitive attributes (sex, race, age) to the outcome variable (income).
We also evaluate all methods on the Neuropathic and Child benchmark networks, averaging results for LLM-based methods across five independent runs.
7.3 Performance Evaluation Metrics
We assess graph reconstruction performance using standard structure-learning metrics measured against the ground-truth DAG.
Precision and Recall.
$\text{Precision}=\frac{\text{TP}}{\text{TP + FP}},\quad\text{Recall}=\frac{\text{TP}}{\text{TP + FN}}$
F1 Score.
[…]
Edge Count Comparison.
We compare the number of predicted edges with the number in the ground-truth graph.
DAG Validation.
Checks acyclicity of the predicted structure.
Normalized Hamming Distance (NHD).
$\text{NHD}=\frac{\text{Mismatches}}{n^{2}}$
where n refers to the number of nodes in the graph.
Adjacency Accuracy.
Accuracy=Correct $\text{Accuracy}=\frac{\text{Correct entries}}{n^{2}}$
8 Results and Analyses
8.1 Causal Discovery Performance Comparison Between Methods
Table 1: Performances of Methods on the Adult-based Synthetic, Child and Neuropathic Causal Networks
Adult-based Synthetic Method Acc. ($\uparrow$) F1 Score ($\uparrow$) Precision Recall NHD ($\downarrow$) NHD Ratio ($\downarrow$) # Pred. Edges
(15 nodes, 28 edges) PC 0.239 0.382 0.352 0.420 0.193 0.743 33
GES 0.296 0.473 0.368 0.580 0.203 0.782 44
NOTEARS ($\lambda=0.01$) 0.021 0.039 0.035 0.045 0.260 0.650 27
DAGMA ($\lambda=0.05$) 0.099 0.180 0.141 0.250 0.283 0.794 50
LLM Pairwise 0.307 0.470 0.331 0.813 0.253 0.530 69
LLM BFS 0.299 0.456 0.332 0.750 0.305 0.539 64
Proposed Method 0.413 0.585 0.792 0.464 0.109 0.415 17
Child PC 0.146 0.255 0.273 0.239 0.097 0.745 22
(20 nodes, 25 edges) GES 0.206 0.341 0.438 0.279 0.119 0.659 16
NOTEARS 0.216 0.356 0.403 0.319 0.080 0.644 20
DAGMA 0.179 0.304 0.333 0.279 0.089 0.696 21
LLM Pairwise 0.130 0.229 0.144 0.559 0.235 0.770 97
LLM BFS 0.150 0.261 0.286 0.240 0.085 0.739 21
Proposed Method 0.364 0.533 0.601 0.479 0.082 0.467 20
Neuropathic PC 0.041 0.078 0.092 0.068 0.025 0.922 563
(221 nodes, 770 edges) NOTEARS 0.022 0.044 0.500 0.023 0.334 0.955 36
DAGMA 0.020 0.039 0.421 0.021 0.351 0.960 38
LLM BFS 0.000 0.000 0.000 0.000 0.903 1.000 43
Proposed Method 0.073 0.136 0.690 0.075 0.109 0.864 84
Table 1 summarizes performance across the synthetic Adult-based, Child, and Neuropathic networks using accuracy, precision, recall, F1 score, and NHD. Across all datasets, the proposed method consistently outperforms baseline approaches in F1 score, precision, and structural alignment with the ground-truth graph. The proposed method generally leads to conservative graph reconstruction compared to the baselines.
Synthetic Adult-based Network.
The proposed method outperforms all baselines in most metrics except recall. It achieves the highest F1 score (0.585), accuracy (0.413), precision (0.792), and the lowest NHD (0.109) and NHD ratio (0.415). This demonstrates its superiority in structural accuracy and edge prediction while producing a more concise graph (17 predicted edges) than other methods.
LLM Pairwise method achieves the highest recall (0.813) but overpredicted edges (69). Classical methods like GES and PC achieved moderate F1 scores (0.473 and 0.382, respectively), but their graphs had higher NHDs and more structural violations. Continuous optimization methods (NOTEARS and DAGMA) underperform across most metrics, indicating sensitivity to noise and difficulty capturing fairness-relevant structure.
Child Network (20 nodes, 25 edges).
The proposed method achieves the highest F1 score (0.533), precision (0.601), and accuracy (0.364), outperforming both statistical and LLM-based baselines. It also achieves the lowest NHD ratio (0.467), indicating better alignment with the ground-truth structure. Although NOTEARS achieves a slightly better raw NHD, our method offers a better precision-recall balance. The LLM Pairwise method shows strong recall (0.559) but overpredicts edges, leading to lower precision and weaker structural alignment.
Neuropathic Network (221 nodes, 770 edges).
The proposed method achieves the highest F1 score (0.136), the highest precision (0.690), and the lowest normalized Hamming distance (NHD) ratio (0.864) across all tested methods. While these metrics are relatively low in absolute terms, they reflect a broader trend where all evaluated methods struggle to recover the full causal structure in this large, high-dimensional graph. This underscores the inherent difficulty of CD in settings characterized by sparse, noisy dependencies and complex variable interactions. Still, the proposed method exhibits higher performance compared to baseline approaches by achieving a better balance between edge prediction accuracy and false positive control.
The LLM-BFS baseline fails entirely on this dataset (F1 = 0, accuracy = 0), likely due to its static query strategy being overwhelmed by the Neuropathic graph’s high dimensionality and sparsity. This underscores the limitations of fixed traversal methods in large-scale settings. In contrast, our method’s use of dynamic scoring and AL improves query prioritization, enabling more effective discovery under complex conditions.
Scalability.
In addition to performance, we also note qualitatively that our method scales more efficiently than optimization-based approaches (e.g., DAGMA, NOTEARS), and the LLM-pairwise method while offering greater structural fidelity. We also perform a complexity analysis described in Appendix E showing how our proposed method could theoretically scale better than the LLM BFS method on significantly large sparse networks, characteristics which real-world networks usually exhibit. This balance that the proposed method provides is evident in the Neuropathic setting, where computationally intensive baselines suffer performance degradation or scalability bottlenecks. Due to these scalability and resource limitations, experiments using the GES and LLM pairwise methods were not completed on the Neuropathic graph.
Reproducibility and Implementation Discrepancies.
We observe lower performance for LLM-BFS and LLM-Pairwise compared to results reported in Jiralerspong et al. (2024), particularly on the Child and Neuropathic networks. This may be due to (1) non-determinism and ongoing updates to models like GPT-4, which may impact consistency across evaluations, especially in semantically ambiguous cases. (2) implementation differences in token limits or message handling, (4) lack of access to original variable descriptions for the Neuropathic dataset, affecting metadata quality. Despite these differences, our implementation adheres to the BFS protocol and reveals its limitations in high-dimensional settings, further motivating our enhancements through dynamic scoring and AL.
8.2 Fairness Pathway Recovery and Bias Attribution
We assess each method’s fairness-awareness by analyzing causal paths from sex, age, and race to income in the Adult-based synthetic network. The ground-truth graph contains 2 direct and 25 indirect fairness paths, with most influence mediated (Indirect = 0.93, Direct = 0.07). A total effect (TE) of 4.89 and normalized bias (C_bias = 28.46) indicate substantial structural bias, aligning with real-world disparities (Appendix D).
We compare each method’s recovered graph in terms of these fairness-aware path metrics. A table detailing the averages of all the fairness and path-related scores for the methods across seed runs is included in Appendix D, as well as a visualization of the graph generated by the outputs of the proposed method.
Figure 2 shows the number of direct, indirect, and spurious paths per method. Due to the wide variance where some methods generate fewer than 5 paths (e.g., NOTEARS), others over 600 (e.g., GES), we use a log-scaled y-axis to preserve readability and comparability without distorting proportions.
Figure 2: Log-scaled path type counts (direct, indirect, spurious) across methods. The log scale enables fair comparison by compressing the large variation in total path counts.
Our proposed method is the only approach to correctly identify both direct fairness paths from sex and race to income, while correctly omitting a spurious direct path from age, demonstrating both sensitivity to true bias and restraint against overgeneralization.
In contrast, LLM Pairwise and BFS methods introduce an incorrect direct path from age, yielding three direct paths in total. This inflated their $C_{\text{bias}}$ values to appear closer to the baseline but was driven by over-attribution rather than accurate recovery. These methods tend to over-attribute causal influence to demographically meaningful variables due to LLM priors and lack of structural constraints, which can amplify perceived bias.
Classical statistical methods, including PC, GES, NOTEARS, and DAGMA, largely underperform. DAGMA and NOTEARS fail to identify any valid paths (direct or indirect) from the sensitive attributes to income, producing exclusively spurious paths. As a result, their $C_{\text{bias}}$ values are near zero, reflecting an inability to detect embedded social bias mechanisms, likely due to their strict dependence on data-driven associations and conservative edge inclusion thresholds.
PC identifies few fairness-relevant paths, and missed both direct edges, leading to a low bias score. This aligns with its constraint-based nature, which can be overly sensitive to noise in independence tests, particularly with sparse or weak signals. GES, while slightly better at recovering indirect effects, also fails to identify direct paths. However, due to its aggressive edge addition during the forward search phase, it captured a larger number of fairness-related pathways overall, majority of which are spurious. This explains its relatively close $C_{\text{bias}}$ score to the baseline, though the accuracy of its causal assignments remains debatable. The graphs constructed by all methods are shown in Appendix D.
Overall, our proposed approach demonstrates a comparatively strong alignment with the fairness-relevant structure of the ground truth. It recovered both true direct effects from sex and race to income, while avoiding spurious paths such as a direct effect from age, which is introduced erroneously by other methods. Although the resulting $C_{\text{bias}}$ value is lower than the baseline, this may reflect the method’s prioritization of high-confidence, high-scoring queries and its dynamic scoring mechanism that discourages over-inclusion. This more conservative behavior potentially reduces the risk of false positives in fairness-sensitive settings, though it may also underestimate total bias when indirect effects are more diffuse. In addition, our method also exhibits the most stable performance across runs, as reflected by its consistently low standard deviation on key metrics such as indirect path count, total effect, and normalized bias as seen in the table in Appendix D. This reliability is important in fairness-sensitive contexts, where reproducibility and robustness are critical.
Figure 3: Comparison of normalized bias ($C_{\text{bias}}$) across methods.
8.3 Parameter Impact Analysis on Proposed Method
We analyze the Bayesian optimization trials using a Random Forest regressor to identify which hyperparameters most influence F1 performance. Max iterations have the greatest impact, followed by score threshold and temperature (Figure 4). This suggests that deeper exploration in AL improves graph construction by reducing false negatives.
Figure 5 supports this, showing F1 scores increasing with iteration budget, up to a plateau, emphasizing the trade-off between computational cost and discovery depth.
Figure 4: Feature importance of hyperparameters in predicting F1 score, computed via Random Forest regression. max_iterations has the highest predictive influence.
Figure 5: Relationship between max_iterations and F1 score. Higher iteration budgets generally yield better causal graph construction performance.
The correlation matrix (Figure 6) shows strong negative correlations among scoring weights, reflecting their competitive role in the dynamic scoring mechanism, while only max iterations correlates positively with F1. This suggests that, for smaller networks like the Adult-based synthetic graph, exploration depth outweighs weight tuning in driving discovery performance.
Figure 6: Pearson correlation heatmap between F1 score and key hyperparameters. Notably, max iterations shows strong positive correlation with F1, while weight terms are moderately anti-correlated.
As graph size increases, parameter importance shifts. In the large Neuropathic network (221 nodes), statistical weights show the strongest positive correlation with F1 (Appendix C), suggesting that MI and PCorr are crucial for narrowing the query space in high-dimensional settings.
Conversely, LLM confidence and query history weights become less influential, likely due to semantic cues being less reliable across deeper or denser graphs. This highlights how our dynamic scoring mechanism could adapt to scale, favoring semantic guidance in small graphs and statistical grounding in larger ones.
9 Discussion
Our findings show that augmenting LLM-based CD with AL and dynamic scoring improves both structural accuracy and fairness sensitivity. The proposed method consistently outperformed baselines across the Adult-based synthetic network, the Child graph, and the large-scale Neuropathic Pain network.
The dynamic scoring mechanism enables effective trade-offs between statistical dependencies and LLM-derived semantics, adapting to graph complexity. This is especially valuable in high-dimensional settings, where statistical cues helped narrow the query space. AL further optimized discovery by prioritizing informative, non-redundant queries.
In terms of fairness, our method is the only one to recover both direct paths from sensitive attributes (sex, race) to income while avoiding false positives (e.g., a spurious link from age). This indicates strong potential for fairness audits with greater precision and reduced over-attribution.
Parameter sensitivity analysis reveals that exploration budget (e.g., max iterations) is the dominant driver of performance. While LLM signals aid early search stages, statistical cues grow in importance with graph size, highlighting the adaptive nature of our hybrid scoring framework across discovery contexts.
While our findings highlight the general potential of CD methods to inform fairness assessments, they also underscore a critical caveat where such methods, particularly when applied naively, can introduce spurious pathways or erroneous connections involving sensitive attributes. Across baseline methods, we observe frequent over or under attribution of causal influence to variables like race, sex, or age, leading to inflated or misleading estimates of structural bias. This is especially problematic in fairness contexts, where false positives (e.g., wrongly inferring a direct path from a sensitive attribute to an outcome), as indicated by the misleadingly similar C_bias scores of graphs from non-performing CD methods to the baseline, may trigger unjustified policy interventions or obscure true sources of inequity. The high variance in graph structure suggests that CD outputs should be interpreted with considerable caution. Practitioners must validate inferred relationships using domain knowledge, robustness checks, and sensitivity analyses before drawing conclusions about discrimination or bias. Without such safeguards, CD risks becoming a source of misinformation rather than a tool for mitigating bias.
10 Limitations
Simplified Causal Modeling in Synthetic Data.
Our synthetic graph, based on the UCI Adult dataset, models income as a terminal node to isolate fairness-related pathways. This design aids in evaluating methods’ ability to detect direct and indirect influences on income. However, it omits downstream effects that income may have in reality—such as on health, wealth, or opportunity, thus excluding feedback dynamics and longitudinal consequences. Future work could extend this framework to capture post-outcome influences and assess fairness in more dynamic causal settings.
Hyperparameter Sensitivity and Computational Constraints.
Although effective, our method requires careful tuning of multiple hyperparameters (e.g., scoring weights, thresholds, LLM temperature), demanding domain expertise. LLM-based querying also incurs substantial computational cost, especially for large networks. Token limits restrict the ability to encode full metadata, particularly in complex graphs like Neuropathic, leading to incomplete or inconsistent responses. Specialized variable names may be misinterpreted, especially at higher temperatures, introducing semantic errors and prediction noise. These factors collectively pose scalability and usability challenges, particularly for resource-constrained settings or high-dimensional domains.
Reproducibility Challenges in LLM-Based Causal Discovery.
Discrepancies between our results and those reported by Jiralerspong et al. (2024) highlight broader reproducibility concerns in LLM-based CD. LLMs are inherently stochastic and sensitive to prompt design, versioning, query order, and API behavior, factors rarely standardized or reported. Additionally, hidden dependencies on tokenization, temperature, or latent state carryover can affect outcomes, making benchmarking difficult. These challenges underscore the need for greater transparency, reproducible prompt templates, and controlled evaluation protocols in future work.
Risk of Social Bias in LLM Responses.
LLMs trained on web-scale data may encode social biases, leading to causal inferences that reflect stereotypes rather than ground-truth mechanisms. This risks semantic hallucinations and biased edge assignments, particularly in fairness-critical applications.
Domain Dependency and Metadata Limitations.
A further limitation lies in the domain dependency of our approach. Because our framework relies on variable metadata to guide LLM-based querying, it is best suited to domains with well-defined, semantically meaningful feature names (e.g., healthcare, census, education). In settings where metadata is sparse, ambiguous, or domain-specific (e.g., sensor networks or high-dimensional genomics), LLM-based judgments may degrade or hallucinate structure (E.g. Neuropathic network). Future work could explore hybrid prompting strategies or domain adaptation techniques to extend semantic CD to less interpretable or evolving data environments.
11 Conclusion
In this work, we propose a fairness-driven CD framework that integrates LLMs with AL and dynamic scoring. Our method outperforms baselines in recovering accurate causal structures across various benchmark networks, and demonstrates more conservative behavior in identifying fairness-relevant pathways, reducing false positives in bias attribution. By combining statistical signals with semantic priors, and prioritizing high-informative queries, the proposed framework advances the reliability and fairness-awareness of LLM-based CD.
To support robust evaluation, we introduce a semi-synthetic benchmark based on the UCI Adult dataset, embedding fairness-relevant causal structure, confounding, and noise. This benchmark provides a controlled yet realistic testbed for future fairness-oriented CD research.
Our hyperparameter analysis revealed that exploration depth (max iterations) strongly influences performance, while the importance of statistical versus LLM-derived signals varies with graph scale.
While our approach improves recovery of fairness-relevant pathways, our framework does not directly translate into end-to-end fairness guarantees in downstream tasks. Bridging the gap between discovery and actionable fairness interventions remains an important area for future integration. Other future directions include expanding to temporal graphs, refining prompt engineering for domain-specific metadata, and building reproducible pipelines for benchmarking LLM-guided causal inference.
Ethical Statement.
This work advances fairness-aware causal discovery by proposing a framework to identify unfair pathways from sensitive attributes to outcomes. The approach can inform the development of more equitable algorithms and policy interventions. However, causal inferences drawn from observational data have inherent limitations and may be misinterpreted without domain expertise. We caution against over-reliance on LLM judgments in high-stakes settings and stress that our method is intended for diagnostic and research purposes only. Additionally, the use of LLMs incurs environmental costs due to their energy-intensive inference. Future work should explore more efficient, bias-aware models to reduce this impact.
Appendix A Appendix: Synthetic Data Generation Rules
This section provides more detail on the Adult-based synthetic data general rules, formulas, and process.
A.1 Data Generation Process
We generate synthetic data based on socioeconomic variables from the UCI Adult dataset. Each variable is sampled according to structural equations encoded in a DAG. Variables are simulated in topological order using a JSON-defined structural causal model with dependency-specific functional forms. The outcome variable income is modeled as a non-linear threshold function over upstream predictors (e.g., capital gain, years of education) and sensitive attributes, mimicking realistic socioeconomic determinants. Data are generated using a Python pipeline with sample sizes of 15,000 and 50,000. Each variable is sampled according to structural equations encoded in a DAG. Table 2 summarizes the variable types, parent dependencies, and sampling formulas included in the baseline graph generation.
Table 2: Variable generation rules for the synthetic dataset (baseline).
Variable Type Dependencies Generation Formula
age Numerical None Uniform(18, 70)
sex Categorical None [”Male”, ”Female”], p = [0.67, 0.33]
race Categorical None [”White”, ”Non white”], p = [0.75, 0.25]
native country Categorical None 5 categories, p = [0.75, 0.1, 0.05, 0.05, 0.05]
education Categorical sex, race, native country Conditional rules on demographics
years of education Numerical education, age $\mathcal{N}(12+2\cdot\text{education}+0.05\cdot\text{age},1)$
occupation Categorical education, sex Rule-based mapping by sex and education
work class Categorical education, race Rule-based mapping by education and race
marital status Categorical age, sex Rule-based by age thresholds and sex
relationship Categorical marital status Rule-based mapping from marital status
capital gain Numerical occupation Normal based on occupation
capital loss Numerical occupation Normal based on occupation
work hours per week Numerical occupation, sex Normal with group-dependent means
fnlwgt Numerical None $\mathcal{N}(180000,20000)$
income Categorical all above Conditional rule combining capital gain, years of education, work hours per week, sex, race
Table 3 configuration describes the synthetic data generation process with noise and latent confounding. All formulas reflect variable dependencies and data-generation logic used in the noisy setup. Here, $U\sim\mathcal{N}(0,1)$ denotes the unobserved latent confounder included in select formulas to simulate realistic social or institutional bias.
Table 3: Noisy Synthetic Data Generation Configuration.
Variable Type Dependencies Formula / Sampling Rule
age Numerical None $\mathcal{U}(18,70)$
sex Categorical (Male, Female) None Probabilities: [0.67, 0.33]
race Categorical (White, Non white) None Probabilities: [0.75, 0.25]
native country Categorical (US, Mexico, India, Philippines, Other) None Probabilities: [0.75, 0.10, 0.05, 0.05, 0.05]
education Categorical (5 levels) sex, race, native country 0 if sex==1 and race==1; 1 if native country $\neq$ 0; 2 if sex==0 and race==0; else 1
years of education Numerical (Discrete) education, age $\mathcal{N}(12+2\cdot\texttt{education}+0.05\cdot\texttt{age}+0.8\cdot U,\ 1)$
occupation Categorical (5 classes) education, sex 2 if education < 1; 1 if sex==0 and education $\geq$ 2; 0 if education $\geq$ 3; else 3
work class Categorical (5 classes) education, race 0 if education > 2; 1 if race==0; 2 if education $\leq$ 1; else 4
marital status Categorical (5 classes) age, sex 0 if age < 30; 1 if sex==0 and age $\geq$ 30; 2 if sex==1 and age $\geq$ 30; else 3
relationship Categorical (4 classes) marital status 0 if marital status==1; 1 if marital status==3; 2 if marital status==0; else 3
[…]
income Categorical (¡=50K, ¿50K) sex, race, education, years of education, occupation, work class, capital gain, capital loss, work hours per week, relationship, marital status, final weight 1 if sex==0 and race==0 and capital gain + 500 $\cdot$ U > 4000 and years of education + 0.5 $\cdot$ U $\geq$ 14 and work hours per week + U > 40; else 0
By explicitly modifying only a few formula components (education, gain, hours, income) and keeping the DAG structure constant, we enable controlled experimentation on how hidden confounding affects CD. This setup supports fairness-sensitive benchmarking while maintaining realism and interpretability.
The baseline synthetic network generated is shown in Figure 7.
Figure 7: Baseline Causal Graph for Adult-based Synthetic Network. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
A.2 Latent Confounding via Unobserved Variable $U$
To reflect unmeasured societal influences (e.g., privilege, institutional access), we introduce a latent confounder $U\sim\mathcal{N}(0,1)$, that affects intermediate variables such as years of education, capital gain, and hours of work per week, thereby indirectly influencing income.
The variables influenced by $U$ were selected based on evidence from the fairness literature and sociological theory. For example, years of education is shaped by early-life educational resources and family background (Nabi and Shpitser 2018; Pearl 2009). Similarly, capital gain reflects socioeconomic access to investments, and work hours are constrained by latent factors like flexibility or family obligations (Nabi and Shpitser 2018; Pearl 2009; Hardt, Price, and Srebro 2016; Heckman 2005).
To simulate these hidden biases, each equation embeds the latent variable $U$ as an additive linear term, following standard structural causal model (SCM) assumptions under additive noise models (ANMs) (Wu et al. 2021; Kilbertus et al. 2017). This models $U$ as an unobserved common cause influencing multiple downstream variables.
Continuous variables are modeled using linear structural equations. Gaussian additive noise models of the form $X_{i}=f(\text{Parents}_{i})+\beta\cdot U_{i}+\epsilon_{i},\ \epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ are used. Here, $X_{i}$ is an observed variable such as years of education or capital gain; $\mu(\cdot)$ represents the deterministic contribution of observed parent variables (e.g., education, age); $\beta\cdot U_{i}$ encodes the latent influence of the unobserved variable $U$; and $\epsilon_{i}$ is a zero-mean residual. For instance, the variable years of education is generated as $\text{Years of Education}\sim\mathcal{N}(12+2\,\text{education}+0.05\,\text{age}+0.8\,U,\ 1)$.
This formulation enables realistic partial confounding, allowing $U$ to simulate latent social advantage while preserving identifiability.It aligns with the LiNGAM framework (Shimizu et al. 2006), structural equation modeling (Pearl 2009), and recent fairness simulation approaches involving latent variables (Kıcıman et al. 2023).
For binary outcomes, such as income, thresholding logic is implemented over both observed and latent inputs using the equation:
$Y=\mathbb{I}\left[f(X)+\lambda U>\theta\right]$ (10)
where $f(X)$ is a (possibly nonlinear) function of observable predictors, $\lambda$ is the strength of the latent effect, and $\theta$ is a threshold. The indicator function $\mathbb{I}[\cdot]$ outputs 1 when the threshold is exceeded. In our setup, this is instantiated as:
[…]
This design captures bias amplification, where latent privilege encoded by $U$ influences outcomes near decision boundaries. Such threshold-based models are commonly used in fairness simulations to mimic opaque real-world decisions (Nabi and Shpitser 2018; Kilbertus et al. 2017; Wu et al. 2021; Kıcıman et al. 2023).
Appendix B Appendix: Experimental Details
In this section, we detail the specific hyperparameters utilized for each set of experiments.
Synthetic Adult-based Network.
We apply six baseline methods (PC, GES, NOTEARS, DAGMA, LLM Pairwise, and LLM BFS) to each dataset instance and averaged performance metrics across seeds. For NOTEARS and DAGMA, we conduct a sweep over multiple regularization values ($\lambda\in\{0.001,0.005,0.01,0.05,0.1\}$) and selected the best performing configuration for reporting: $\lambda=0.01$ for NOTEARS and $\lambda=0.05$ for DAGMA.
We perform 200 trials per dataset, optimizing for F1 score via Bayesian search. The tuned hyperparameter space includes: query prioritization weights selected from $[(0.3,0.3,0.4),(0.25,0.25,0.5),(0.4,0.3,0.3),(0.3,0.4,0.3)]$, score threshold values between 0.25 and 0.35, temperature values between 0.25 and 0.6, and a maximum iteration range of 40 to 100. We report the average across seeds for all performance metrics used in evaluation.
Fairness pathway analysis compares recovered and true graphs to assess whether each method correctly identifies bias mechanisms. CD outputs include (1) Causal Graphs: Represented by adjacency matrices capturing the graph structure, and (2) Causal Dictionaries: Capturing identified causal relationships. We focus on paths from sensitive attributes (sex, race, age) to the outcome variable (income) for fairness evaluation.
Child Network.
We perform 1000 trials across all chunks for the proposed method. Evaluation is based on the F1 score with respect to the ground-truth DAG. Hyperparameter ranges are: score threshold $[0.01,0.2]$, temperature $[0.05,0.7]$, and max iterations $[10,50]$.
Neuropathic Pain Network.
We run 200 extended trials for the proposed method to explore effects on long-range dependencies. Parameter ranges include score thresholds $[0.1,0.5]$, temperature $[0.1,0.3]$, and max iterations $[200,500]$.
Appendix C Appendix: Parameter Impact Analysis on Proposed Method For Child and Neuropathic Networks
To evaluate the influence of various parameters on the performance of the proposed method, a detailed analysis of the hyperparameters is conducted using visualizations such as ridge plots, correlation heatmaps, and feature importance analysis.
Impact of Weights on Performance.
Figure 8 shows ridge plots illustrating the impact of the 3-weight combination on the F1 score for experiments on the Child network. Each plot visualizes F1 score distributions across different weight bins. The width indicates variation in F1 scores; a wider ridge means more variation, while a narrower ridge suggests consistent performance. The height of the peak shows the most frequently occurring F1 scores, with a taller peak indicating a common F1 score for that weight bin. Colors distinguish ridges for each bin without indicating performance or significance. The figure shows that statistical information and query history are significant contributors to performance, while LLM confidence has a stable, less pronounced effect.
  • Statistical Information Weight: Lower bins (0.0-0.1) show narrower ridges and lower F1 scores. As the weight increases (0.3-0.6), distributions broaden, and peaks shift toward higher F1 scores, suggesting optimal performance in this range. Beyond 0.6, performance flattens with diminishing returns.
  • LLM Confidence Weight: Distributions remain stable across all bins, with minor improvements in midrange bins (0.3-0.5). This weight has a minor impact on the F1 score and serves as a supporting factor rather than a decisive one.
  • Query History Weight: Lower bins (0.0-0.2) show lower F1 scores. As the weight increases (0.4-0.6), distributions widen, and peaks shift toward higher F1 scores. Beyond 0.6, improvement plateaus, with optimal performance around 0.4-0.6.
These results emphasize the importance of balancing weights, as increasing one decreases the others due to the constraint that all weights must sum to 1.
Figure 8: Ridge Plots for Weights of Statistical Information, LLM Confidence, and Query History against F1-Score
Figure 9: Correlation Heatmap of Parameters and F1 Score - Neuropathic Network
Correlation Analysis.
Figure 9 illustrates the relationships between model parameters and the F1 score, revealing key performance factors for experiments from the neuropathic dataset. The weight of statistical information has the strongest positive correlation with the F1 score (+0.27), indicating that increasing its value generally improves performance. The weight of the query history shows a weaker negative correlation (-0.17), reflecting diminishing returns beyond its optimal range (0.4-0.6). The weight of the LLM confidence score has a minimal impact (-0.09 correlation), therefore, its optimization is less critical. Negative correlations between weights (e.g., -0.54 between statistical information and query history) highlight trade-offs due to their sum equaling 1. The score threshold shows negligible influence on the F1 score and can be deprioritized. Temperature and the number of maximum iterations have weak positive correlations (+0.02 and +0.18, respectively), so small adjustments have limited impact.
These findings underscore the importance of optimizing statistical information and query history weights while treating other parameters as secondary contributors to model performance.
Appendix D Appendix: Fairness and Path Analyses Results
This table shows the average fairness and path scores for the baseline synthetic graph and the graphs generated by the CD methods.
Table 4: Fairness and Causal Pathway Analysis across Methods (Averaged across 4 seeds)
Method Direct Indirect Spurious Fairness Path Direct Effect Indirect Effect Total Effect Normalized C_Bias
Paths Paths Paths Contribution
Baseline (GT) 2 25 25 0.52 0.623 4.272 4.894 28.465
Proposed Method 2 3 $\pm$ 1 5 $\pm$ 1 0.467 $\pm$ 0.021 0.519 0.247 $\pm$ 0.003 0.766 $\pm$ 0.002 4.602 $\pm$ 0.016
LLM BFS 3 44 $\pm$ 18 121 $\pm$ 10 0.273 $\pm$ 0.061 0.519 2.930 $\pm$ 0.47 3.447 $\pm$ 0.46 20.699 $\pm$ 2.58
LLM Pairwise 3 222 $\pm$ 99 281 $\pm$ 114 0.443 $\pm$ 0.05 0.519 6.136 $\pm$ 2.76 6.655 $\pm$ 2.66 39.992 $\pm$ 15.4
NOTEARS (0.01) 0 0 5 $\pm$ 6 0.00 0.00 0.00 0.00 0.00
DAGMA (0.05) 0 0 31 $\pm$ 2 0.00 0.00 0.00 0.00 0.00
GES 0 45 $\pm$ 10 638 $\pm$ 100 0.068 $\pm$ 0.03 $\sim$0 2.892 $\pm$ 0.32 6.293 $\pm$ 1.81 13.975 $\pm$ 7.35
PC 0 1 $\pm$ 1 51 $\pm$ 23 0.02 $\pm$ 0.01 0.00 0.221 $\pm$ 0.12 0.221 $\pm$ 0.12 1.326 $\pm$ 1.05
The graphs generated using the proposed method, PC, GES, DAGMA, Notears, LLM Pairwise, and LLM BFS methods are shown in Figures 10 to 16.
Figure 10: Causal Graph for Adult-based Synthetic Network Generated by Proposed Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
Figure 11: Causal Graph for Adult-based Synthetic Network Generated by PC Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
Figure 12: Causal Graph for Adult-based Synthetic Network Generated by GES Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
Figure 13: Causal Graph for Adult-based Synthetic Network Generated by Dagma Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
Figure 14: Causal Graph for Adult-based Synthetic Network Generated by Notears Method. Red and orange nodes represent target and sensitive attributes respectively, blue nodes represent others.
Figure 15: Causal Graph for Adult-based Synthetic Network Generated by LLM Pairwise Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
Figure 16: Causal Graph for Adult-based Synthetic Network Generated by LLM BFS Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
Appendix E Appendix: Complexity Analyses
The original BFS method (Jiralerspong et al. 2024) exhaustively explores variable relationships without prioritization, resulting in high computational complexity. It queries independent nodes with complexity $\mathcal{O}(n)$, where $n$ is the number of variables. For each variable $x$, the LLM is queried about all potential children $y$, leading to a total query complexity of $\mathcal{O}(n^{2})$. Cycle prevention has complexity $\mathcal{O}(n+e)$, where $e$ is the number of edges, resulting in a worst-case overall complexity of $\mathcal{O}(n^{3})$.
The proposed method introduces AL and a Dynamic Scoring Mechanism to reduce the number of queries. Querying independent nodes remains $\mathcal{O}(n)$. Pre-computing scores for all pairs $(x,y)$ has complexity $\mathcal{O}(n^{2})$. Dynamic scoring during each iteration involves updating scores, with complexity $\mathcal{O}(k\cdot(n^{2}-q))$, where $q$ is the number of queried pairs and $k$ is the maximum number of iterations. This approach prioritizes high-scoring pairs, reducing $q$ compared to the original BFS method. Cycle prevention complexity is $\mathcal{O}(q^{\prime}\cdot(n+e))$, where $q^{\prime}\ll n^{2}$. The overall complexity is $\mathcal{O}(n^{2}+k\cdot q^{\prime}+q^{\prime}\cdot(n+e))$. This significantly improves efficiency over the original BFS approach. Table 5 summarizes these improvements.
Definition of $q$ and $q^{\prime}$. We define:
  • $q$: The number of variable pairs actually queried by the LLM.
  • $q^{\prime}=n^{2}-q$: The number of unqueried or skipped variable pairs due to low informativeness.
The scoring function is updated only for a subset of $q^{\prime}$, typically involving nodes affected by newly added edges, thus avoiding global score recomputation.
Table 5: Computational Complexities of Proposed Method Vs. BFS Method
Step BFS method (Jiralerspong et al. 2024) Proposed Method Improvement
Independent Nodes $O(n)$ $O(n)$ No improvement
Querying Relationships $O(n^{2})$ $O(k\cdot q^{\prime})$ Reduces queries through prioritization.
Dynamic Scoring Not applicable $O(n^{2}+k\cdot q^{\prime})$ Prioritizes high-impact queries.
Cycle Detection $O(n^{2}\cdot(n+e))$ $O(q^{\prime}\cdot(n+e))$ Fewer edges evaluated due to early stopping.
Overall Complexity $O(n^{3})$ $O(n^{2}+k\cdot q^{\prime})$ Significant reduction in queries.

### v2

@@
 Abstract
-Causal discovery (CD) plays a pivotal role in understanding the mechanisms underlying complex systems.
-While recent algorithms can detect spurious associations and latent confounding, many struggle to recover fairness-relevant pathways in realistic, noisy settings.
-Large Language Models (LLMs), with their access to broad semantic knowledge, offer a promising complement to statistical CD approaches, particularly in domains where metadata provides meaningful relational cues.
-Ensuring fairness in machine learning requires understanding how sensitive attributes causally influence outcomes, yet CD methods often introduce spurious or biased pathways.
-We propose a hybrid LLM-based framework for CD that extends a breadth-first search (BFS) strategy with active learning and dynamic scoring.
-Variable pairs are prioritized for LLM-based querying using a composite score based on mutual information, partial correlation, and LLM confidence, improving discovery efficiency and robustness.
-To evaluate fairness sensitivity, we construct a semi-synthetic benchmark from the UCI Adult dataset, embedding a domain-informed causal graph with injected noise, label corruption, and latent confounding.
-We assess how well CD methods recover both global structure and fairness-critical paths (e.g., sex $\rightarrow$ education $\rightarrow$ income).
-Our results show that LLM-guided methods, including the proposed method, demonstrate competitive or superior performance in recovering such pathways under noisy conditions.
-We highlight when dynamic scoring and active querying are most beneficial and discuss implications for fairness auditing in real-world datasets.
+Ensuring fairness in machine learning requires understanding how sensitive attributes like race or gender causally influence outcomes.
+Existing causal discovery (CD) methods often struggle to recover fairness-relevant pathways in the presence of noise, confounding, or data corruption.
+Large language models (LLMs) offer a complementary signal by leveraging semantic priors from variable metadata.
+We propose a hybrid LLM-guided CD framework that extends a breadth-first search strategy with active learning and dynamic scoring.
+Variable pairs are prioritized for querying using a composite score combining mutual information, partial correlation, and LLM confidence, enabling more efficient and robust structure discovery.
+To evaluate fairness sensitivity, we introduce a semi-synthetic benchmark based on the UCI Adult dataset, embedding domain-informed bias pathways alongside noise and latent confounders.
+We assess how well CD methods recover both global graph structure and fairness-critical paths (e.g., sex→education→income).
+Our results demonstrate that LLM-guided methods, including our active, dynamically scored variant, outperform baselines in recovering fairness-relevant structure under noisy conditions.
+We analyze when LLM-driven insights complement statistical dependencies and discuss implications for fairness auditing in high-stakes domains.
 1 Introduction
-Causal discovery (CD) identifies causal relationships among variables and is widely used in domains such as epidemiology, genetics, and economics (Kampani et al.
+Bias in machine learning (ML) systems affects decisions in hiring, lending, education, and healthcare.
+These biases often emerge through indirect, structural pathways where sensitive attributes (e.g., race or gender) influence outcomes via proxies or confounded relationships (Graetz, Boen, and Esposito 2022).
+Standard fairness audits rely on statistical disparity metrics, but such metrics overlook how bias propagates causally (Chinta et al.
+2025).
+Without structural insight, interventions may be ineffective or misleading.
+Causal discovery (CD) offers tools to identify such pathways, distinguishing genuine effects from those introduced by confounders or measurement artifacts.
+While recent methods apply structural causal models or fairness constraints to audit ML systems, classical CD techniques often fail under noisy data, latent confounding, or incomplete metadata.
+These limitations are acute in fairness contexts where both precision and interpretability are critical (Takayama et al.
 2024).
-Recent advances in CD algorithms have opened new avenues for addressing bias and fairness in machine learning (ML).
-Bias in ML often arises from hidden confounders or unequal treatment of sensitive groups (Barocas, Hardt, and Narayanan 2023).
-For instance, a hiring algorithm may inherit biased patterns from historical data that unfairly disadvantage certain demographic groups (Binns 2018).
-CD helps identify such pathways, distinguishing genuine effects from biases introduced by confounders and ensuring decisions are based on causal factors rather than proxies (Makhlouf, Zhioua, and Palamidessi 2020).
-Unlike traditional methods relying on statistical correlations or predefined fairness metrics, CD disentangles direct effects, indirect effects, and spurious correlations (Miguel-Hurtado et al.
-2016).
-This has prompted growing interest in using CD to support fairness-aware ML.
-Recent studies have proposed using counterfactual reasoning, structural causal models (SCMs), and fairness-specific constraints to detect and mitigate discriminatory pathways (Binkytė et al.
-2023; Zuo et al.
-2022; Wang and Luo 2023).
-However, classical CD methods often rely on strong assumptions, perform poorly under noise or confounding, and struggle to align with domain knowledge not explicitly encoded in data (Takayama et al.
+Large language models (LLMs) have emerged as promising tools for CD, using their vast semantic knowledge to infer causal directions between variables from textual metadata (Le, Xia, and Chen 2024; Vashishtha et al.
+2023).
+Prior work has used LLMs to infer causal links or orderings, and to reduce query cost via strategies like breadth-first search (BFS) (Jiralerspong et al.
 2024).
-Large Language Models (LLMs) have emerged as promising tools for CD, using their vast semantic knowledge to infer causal directions between variables such as “sex” and “income” from textual metadata (Le, Xia, and Chen 2024; Vashishtha et al.
-2023; Khatibi et al.
-2024; Kampani et al.
-2024; Long, Schuster, and Piché 2023; Kıcıman et al.
-2023).
-Jiralerspong et al.
-(Jiralerspong et al.
-2024) proposed a breadth-first search (BFS) strategy for efficient graph construction via LLM queries, showing improved performance over random or exhaustive querying.
-However, a growing concern is that CD methods, especially when used in fairness audits, may introduce or amplify bias by inferring spurious links involving sensitive features.
-In fairness-sensitive contexts, incorrect causal attributions can lead to unjustified conclusions about discrimination or fairness violations.
-Thus, ensuring both the accuracy and interpretability of discovered pathways is essential for responsible ML deployment.
-Building on prior work, we introduce an enhanced BFS framework that integrates Active Learning (AL) and Dynamic Scoring.
-Instead of querying all pairs uniformly, our method ranks pairs using mutual information (MI), partial correlation (PCorr), and LLM-derived confidence, while discouraging redundant queries and focusing on informative regions via a history-based weight.
-This approach treats the LLM not as a passive responder but as an adaptive component within an iterative learning loop.
+However, naive LLM use risks over or under attributing causality, especially involving sensitive attributes, raising concerns about spurious fairness conclusions.
+Building on prior work, we introduce a fairness-driven CD framework that enhances the BFS approach with active learning (AL) and dynamic scoring.
+Our method prioritizes variable pairs using mutual information (MI), partial correlation (PCorr), and LLM-derived confidence, while discounting redundant queries and focusing on informative regions via a history-based weight.
+Unlike exhaustive querying or fixed-order search, our adaptive strategy balances semantic and statistical cues to recover fairness-relevant paths more efficiently.
 A key challenge in evaluating CD methods in fairness contexts is the absence of real-world datasets with known ground-truth causal structures involving sensitive attributes (Loftus et al.
@@
 Most, like the UCI Adult dataset, offer rich features but lack verified causal graphs.
-This limitation makes it difficult to validate whether inferred pathways truly reflect bias mechanisms or are artifacts of the discovery process.
-To address this, we construct a semi-synthetic benchmark that retains real-world semantics while embedding a known fairness-relevant graph for controlled evaluation.
-We analyze the performance of our proposed method and other baseline CD methods on the constructed semi-synthetic benchmark based on the UCI Adult dataset that incorporates fairness-relevant structure, noise, and latent confounding.
-We evaluate these methods across their ability to recover key pathways (e.g., sex $\rightarrow$ education $\rightarrow$ income) and to support fairness audits.
+To enable robust evaluation, we construct a semi-synthetic benchmark from the UCI Adult dataset, embedding a known fairness-relevant causal graph with noise, corruption, and confounding.
+We assess how well CD methods recover both global structure and key fairness pathways (e.g., sex $\rightarrow$ education $\rightarrow$ income).
 We also conduct a hyperparameter sensitivity analysis to examine how LLM-specific factors such as temperature, query prioritization weights, and stopping criteria affect CD performance.
 These findings offer practical insights into how LLM behavior and prompting influence graph reconstruction accuracy and fairness-path recovery.
-Our contributions are as follows:
-  • We propose a hybrid LLM-based CD method that integrates AL and dynamic scoring to prioritize informative queries during BFS-based graph recovery.
-  • We introduce a semi-synthetic benchmark grounded in fairness literature, with latent confounding, label noise, and sensitive pathways, to evaluate CD methods under realistic conditions.
-  • We provide a comparative analysis of pathway recovery across methods and examine how different discovery strategies affect fairness-related inferences.
-  • We analyze the impact of LLM-specific hyperparameters on discovery quality.
+Our approach is designed for robustness, interpretability, and social accountability, enabling stakeholders to trace how outcomes are structurally linked to sensitive variables and to intervene in meaningful ways.
+Our key contributions:
+  • A hybrid LLM-CD framework with active learning and dynamic scoring to enhance fairness-aware structure discovery.
+  • A benchmark for fairness-sensitive CD grounded in real-world semantics and controlled ground truth.
+  • An evaluation of pathway recovery across methods, highlighting conditions where LLM-guided discovery improves fairness diagnostics.
+  • A sensitivity analysis on LLM parameters and their influence on recovery quality.
 2 Related Work
-Causal discovery has been explored through statistical, optimization-based, and more recently, LLM-guided approaches (Long et al. 2025).
-Classical methods like the Peter-Clark (PC) algorithm (Spirtes and Glymour 1991) and Greedy Equivalence Search (GES) (Meek 1997) use conditional independence tests or score-based search but rely on strong assumptions (e.g., causal sufficiency, faithfulness) and degrade with latent confounding or limited data.
-Optimization-based approaches, such as NOTEARS (Zheng et al.
-2018) and DAGMA (Bello, Aragam, and Ravikumar 2022), frame CD as a constrained optimization problem.
-While accurate, they are computationally expensive and sensitive to hyperparameters, limiting their use in fairness-critical or iterative settings.
-LLM-based CD methods offer an alternative by leveraging variable descriptions to infer structure, offering a way to encode human knowledge and priors.
-Kıcıman et al.
+Statistical and Optimization-Based CD.
+CD methods span statistical, optimization-based, and more recently, LLM-guided approaches (Long et al.
+2025).
+Classical techniques such as PC (Spirtes and Glymour 1991) and GES (Meek 1997) rely on conditional independence tests and strong assumptions (e.g., faithfulness), often failing under latent confounding or limited data.
+Optimization-based methods like NOTEARS (Zheng et al.
+2018) and DAGMA (Bello, Aragam, and Ravikumar 2022) offer greater precision but are computationally expensive and sensitive to hyperparameters, making them less practical for fairness-critical contexts.
+LLM-Guided CD.
+Recent LLM-based approaches use variable semantics to infer structure, with Kıcıman et al.
 (Kıcıman et al.
-2023) introduced pairwise LLM queries using only metadata, though the quadratic query complexity limits scalability.
-Others combine LLM-derived causal orders with algorithms like PC (Vashishtha et al.
-2023), or embed LLMs in multi-agent or prompt-refinement loops (Khatibi et al.
+2023) proposing pairwise causal queries from metadata.
+However, such methods scale poorly due to quadratic query complexity.
+Others combine LLM-inferred causal orders with classical CD (Vashishtha et al.
+2023), or use multi-agent prompting strategies (Khatibi et al.
 2024; Le, Xia, and Chen 2024).
@@
 (Kampani et al.
-2024) take a hybrid approach, using LLMs to generate priors, refining them with methods like NOTEARS.
+2024) generate LLM-based priors refined through NOTEARS.
 Takayama et al.
 (Takayama et al.
-2024) propose statistical causal prompting, where LLMs enhance statistical methods by providing priors or plausible causal orders.
-While robust, this one-time augmentation lacks iterative refinement.
-Similarly, Jiralerspong et al.
+2024) introduced statistical causal prompting, where LLMs enhance statistical methods by providing priors or plausible causal orders.
+Jiralerspong et al.
 (Jiralerspong et al.
-2024) propose a BFS framework to reduce LLM query complexity during structure discovery.
-Their approach treats LLM responses as inputs for efficient edge discovery, dramatically lowering runtime without sacrificing metadata grounding.
-However, BFS applies fixed selection heuristics and does not dynamically balance statistical evidence with LLM insights, making it susceptible to early decision errors.
-Our method builds on this BFS foundation, introducing AL and a dynamic scoring strategy that prioritizes variable pairs based on a mixture of MI, PCorr, and LLM confidence scores.
-Compared to prior methods that treat all variable pairs uniformly or rely solely on LLMs or statistics, our approach dynamically balances complementary signals and avoids over-reliance on any single information source.
-Some of the above-mentioned approaches differ fundamentally from our framework in both scope and evaluation goals.
-Methods such as those proposed by Vashishtha et al.
-(Vashishtha et al.
-2023), Khatibi et al.
-(Khatibi et al.
-2024), Le et al.
-(Le, Xia, and Chen 2024), and Kampani et al.
-(Kampani et al.
-2024) use LLMs to generate global causal orders or priors that are later refined using constraint-based or optimization-based algorithms like PC or NOTEARS.
-In contrast, our method treats the LLM as an active querying agent embedded in an iterative loop, where causal judgments are elicited at the pairwise level and guided dynamically by a scoring function informed by both statistical and historical signals.
-Unlike prompting-based or prior-initialized approaches, our framework does not rely on a static, up-front extraction of priors, but instead continuously adapts query selection through AL.
-We further emphasize fairness-aware causal recovery, targeting sensitive paths and mediators (e.g., sex $\rightarrow$ education $\rightarrow$ income).
-This responds to calls for causal approaches to fairness (Nabi and Shpitser 2018; Kilbertus et al.
-2017; Zhang and Bareinboim 2018; Chiappa 2019), highlighting the need for explainable, intervenable, and robust fairness assessments under noisy conditions.
-In this sense, our method contributes to a growing understanding of how LLM-based querying, when combined with statistical reasoning, can affect fairness-relevant causal inference under noisy and complex conditions.
+2024) proposed a BFS strategy to reduce query complexity, but rely on fixed pair ordering, making them susceptible to early decision errors.
+We build on this line by introducing AL and a dynamic scoring mechanism to adaptively prioritize informative variable pairs using MI, PCorr, and LLM confidence.
+Our method dynamically balances semantic and statistical signals, in contrast to prior work that uses LLMs for one-time priors or uniform querying.
+We focus on observational CD methods to maintain comparability with LLM-based approaches, leaving integration with interventional or experimental methods (e.g., do-calculus, IV-based CD) for future work.
+Fairness and Causal Pathways.
+Beyond structural accuracy, recent work has emphasized the importance of CD for fairness auditing.
+Causal fairness frameworks (Nabi and Shpitser 2018; Kilbertus et al.
+2017; Zhang and Bareinboim 2018; Chiappa 2019) focus on identifying discriminatory pathways, e.g., sex $\rightarrow$ education $\rightarrow$ income that explain how bias propagates through mediation or proxy variables.
+Our method is explicitly tuned to recover such fairness-relevant paths with greater robustness under noise and limited samples, enabling more actionable fairness assessments.
+Social and Regulatory Context.
+Beyond computer science, our work connects to social science perspectives.
+Structural discrimination and intersectionality highlight how compounding disadvantage flows through proxy variables, not just direct links (Crenshaw 2013; Bonilla-Silva 1997).
+Regulatory frameworks such as GDPR’s “right to explanation” and EEOC compliance requirements also motivate interpretable, path-level audits (Kaminski 2021).
+Causal Inference and Accountability.
+We also draw from causal inference frameworks like the potential outcomes model (Rubin 1974; Imbens and Rubin 2015) and matching-based fairness analysis (Stuart 2010), which stress structural validity and transparency of assumptions.
+Our hybrid design reflects these values by surfacing interpretable causal hypotheses grounded in both observed data and language-based priors.
+Finally, we respond to calls for auditable AI systems through transparency-centered work on model documentation and sociotechnical accountability (Mitchell et al.
+2019; Selbst et al.
+2019).
 3 Proposed LLM-Based Method
-3.1 LLM-Guided Breadth-First Search for Causal Discovery
-The foundation of the proposed method is the BFS-based CD framework introduced by (Jiralerspong et al.
-2024).
-This approach iteratively queries an LLM to infer causal relationships among variables in a data set and reduces the number of queries required since visiting each node once is sufficient to determine all edges originating from that node.
-Using this prompting approach, each query is approached as an opportunity for node expansion within a BFS algorithm, progressively building the causal graph as it is traversed using BFS.
-To effectively implement the BFS, it is essential to determine the sequence in which nodes are explored.
-For this purpose, the Directed Acyclic Graph (DAG) nature of causal graphs is leveraged, by using their topological ordering to guide the BFS traversal.
-The BFS process follows three main steps:
-  1. 1.
-    Initialization: The algorithm begins by querying the LLM to identify independent variables that are not caused by any other variables which form the starting points for graph exploration.
-  2. 2.
-    Expansion: For each variable on the current frontier (independent variables or those already explored), the LLM is queried about which other variables it causally influences, using prompts that incorporate discovered structure so far.
-  3. 3.
-    Graph Construction: Inferred edges are added to a directed graph if they do not introduce cycles.
-Each new relationship is checked to ensure that the variables in the frontier are connected to previously unexplored nodes.
-The process continues until all nodes are explored.
-The final output is a predicted causal graph in the form of an adjacency matrix.
-While more efficient than exhaustive pairwise querying, this approach still explores variable pairs uniformly.
-As a result, many LLM queries target uninformative pairs, especially in sparse or noisy graphs, motivating the need for adaptive query selection.
+We build on the BFS-based CD framework of (Jiralerspong et al.
+2024), where an LLM is queried iteratively to construct a causal graph.
+The method begins by identifying independent root variables and expands outward via LLM-guided queries, treating each confirmed edge as a BFS traversal step.
+To ensure acyclicity, edges are added only if they do not create cycles.
+While more efficient than exhaustive querying, this approach treats all variable pairs uniformly and can waste queries on uninformative or redundant relationships.
 3.1 Active Learning and Dynamic Scoring for Efficient Querying
-To improve the informativeness of queries and reduce redundancy, we introduce a scoring-based AL framework.
-This component augments the BFS approach with a dynamic prioritization mechanism for selecting the most valuable variable pairs to query.
-Figure 1 illustrates the full framework.
+We introduce an AL strategy that selects variable pairs based on a dynamic informativeness score.
+This score integrates statistical signals, model confidence, and query history, prioritizing pairs likely to yield informative causal judgments.
+The query loop continues until a maximum iteration or informativeness threshold is reached.
+Figure 1 summarizes the overall architecture.
 Figure 1: Overview of the proposed LLM-guided BFS framework with dynamic scoring and AL.
 Dynamic Scoring Mechanism
-For each unqueried pair $(x,y)$, a dynamic informativeness score $S(x,y)$ is computed as:
+Each unqueried pair $(x,y)$ receives a composite score:
 $\begin{split}S(x,y)=\ &w_{\text{stat}}\cdot\text{StatScore}(x,y)+\\ &w_{\text{conf}}\cdot\text{LLMConf}(x,y)+w_{\text{hist}}\cdot\text{HistScore}(x,y)\end{split}$ (1)
 Each term incorporates a distinct signal:
-  • Statistical Score: Combines mutual information (MI) and partial correlation (PCorr), measuring both linear and non-linear dependencies:
+  • Statistical Score: Combines MI and PCorr, measuring both linear and non-linear dependencies:
     $\text{StatScore}(x,y)=\frac{\text{MI}(x,y)+\text{PCorr}(x,y)}{2}$ (2)
+    PCorr is computed conditionally on the current discovered parent sets, following the iterative conditioning procedure of the PC algorithm.
   • LLM Confidence Score: Reflects the certainty of the model’s previous response for a given pair, where higher token-level confidence leads to higher scores:
@@
 These weights are optimized via Bayesian optimization (see Section 6), allowing the framework to adapt scoring behavior to different graph structures or domain characteristics.
-Query Selection via Active Learning
-We incorporate dynamic scoring into an AL loop that determines which pairs to query. The query process iteratively selects:
-$(x^{*},y^{*})=\arg\max_{(x,y)\in\text{UnqueriedPairs}}S(x,y)$ (5)
-An LLM is then prompted to assess whether $x^{*}$ causally affects $y^{*}$ using the natural language format described in listing 1.
-Listing 1 Causal query prompt format
-Does X cause Y? Provide the answer in the format: <Answer>Yes</Answer> or <Answer>No</Answer>.
-If the LLM predicts a positive causal link and the resulting graph remains acyclic, the edge $x^{*}\rightarrow y^{*}$ is added; otherwise, it is discarded.
+Querying the LLM
+At each step, the pair $(x^{*},y^{*})$ with the highest score is selected:
+$(x^{*},y^{*})=\arg\max_{(x,y)\in\text{Unqueried}}S(x,y)$ (5)
+The LLM is queried using a simple natural language prompt:
+Listing 1 Simplified LLM Query Format
+1System: You are a domain expert.
+2User: You assist the expert in evaluating possible causal links.
+3
+4Metadata: Each variable is described using a short name and a natural language description.
+5
+6Query: Does VAR_A cause VAR_B?
+7Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
+Domain-specific instructions and variable descriptions are included to contextualize each query. See Appendix for full example prompts.
+If the LLM returns “Yes” and adding $x^{*}\rightarrow y^{*}$ does not form a cycle, the edge is added to the graph.
 To estimate the LLM uncertainty, we extract token-level log-probabilities and compute the average probability of the model’s response (“Yes” or “No”).
@@
 The process terminates after a query limit or when scores fall below a threshold.
-Cycle Detection and Output Construction
 To ensure the output is a valid DAG, each edge addition is immediately followed by a cycle check. If the graph becomes cyclic, the addition is reversed.
@@
 The entire procedure is conducted in a multi-turn chat session, allowing the LLM to maintain contextual awareness of previously inferred relationships and rationales.
-4 Fairness Evaluation through Causal Pathway Analysis
-To evaluate the fairness utility of learned causal graphs, we perform pathway analysis to quantify how sensitive variables influence target outcomes.
-Unlike classical fairness metrics based on statistical associations (e.g., demographic parity), pathway analysis distinguishes between direct, mediated, and spurious effects, offering a principled causal perspective to fairness assessment (Pearl 2009, 2022).
-We analyze the structure and influence of paths from sensitive attributes $S$ (e.g., race, sex) to outcome variables $Y$ (e.g., income) in the learned graphs.
-This approach distinguishes between direct, mediated, and spurious influences, going beyond traditional statistical fairness metrics that lack causal grounding (Pearl 2022).
-Path Count Comparison.
-For each discovered graph, we enumerate and classify all directed paths originating from sensitive variables:
-  • Direct Paths ($P_{\text{direct, fairness}}$): Number of direct causal edges ($S\rightarrow Y$) that connect the defined sensitive variables to the outcomes.
-  • Indirect Paths ($P_{\text{indirect, fairness}}$): Number of causal paths ($S\xrightarrow{}\text{intermediate nodes}\xrightarrow{}Y$) in which sensitive variables influence the outcomes indirectly via intermediate variables.
-  • Spurious Paths ($P_{\text{spurious, fairness}}$): Number of paths that involve sensitive variables but do not influence the outcomes ($S\xrightarrow{}\text{non-$Y$ nodes}$).
-To assess structural alignment with fairness-relevant pathways, we report the number of direct and indirect paths for each graph and compare these across CD methods and to the known ground-truth graph.
-A higher overlap with true fairness-relevant paths indicates better recovery of bias mechanisms.
-Quantitative Effect Decomposition.
-To assess the magnitude of fairness-related effects, we estimate:
-  • Direct Effect (DE): The estimated causal effect of $S$ on $Y$ along direct paths.
-  • Indirect Effect (IE): The effect of $S$ on $Y$ through mediating variables.
-  • Total Effect (TE): The sum of direct and indirect effects:
+4 Fairness Evaluation via Pathway Analysis
+We assess the fairness utility of learned causal graphs by analyzing how sensitive attributes $S$ (e.g., race, sex) influence outcomes $Y$ (e.g., income).
+Unlike statistical fairness metrics, pathway analysis provides a causal perspective by distinguishing direct, mediated, and spurious effects (Pearl 2009, 2022).
+Path Classification
+We enumerate all directed paths from $S$ and classify them as:
+  • Direct: Edges $S\rightarrow Y$.
+  • Indirect: Paths $S\rightarrow\cdots\rightarrow Y$ via mediators.
+  • Spurious: Paths involving $S$ that do not reach $Y$.
+Comparing these across methods and against the ground-truth graph reveals how well fairness-relevant mechanisms are recovered.
+Effect Decomposition
+We estimate the causal effects of $S$ on $Y$ using either structural equations or interventional estimators:
+$DE$ $:\text{Direct effect (via }S\rightarrow Y\text{)}$
+$IE$ $:\text{Indirect effect (via mediators)}$
 $TE$ $=DE+IE$
-These effects are computed using structural equations or interventional estimators (depending on the graph), and normalized to account for scale:
+We normalize by outcome variance to obtain:
 $C_{\text{bias}}=\frac{TE}{\text{Var}(Y)}$ (7)
-where $C_{\text{bias}}$ represents the total fairness-relevant contribution of $S$ to variance in the outcome.
-This enables comparative evaluation across datasets and methods.
-It quantifies the total effect of sensitive attributes on the outcome (income), scaled by outcome variance.
-A higher $C_{\text{bias}}$ indicates a greater risk of propagating sensitive information through direct or mediated pathways, particularly important for fairness auditing.
-5 Synthetic Graph Construction Based on UCI Adult Attributes
-To evaluate fairness-aware CD under controlled yet realistic conditions, we construct a semi-synthetic dataset based on the UCI Adult dataset (Kohavi et al.
-1996), commonly used in fairness research for its demographic diversity and income disparities.
-Our synthetic graph retains the original 15 variables and encodes plausible causal links informed by prior work (Kilbertus et al.
+Here, $C_{\text{bias}}$ quantifies the fairness-relevant contribution of $S$ to $Y$, enabling method comparison across datasets. Higher values indicate a greater risk of bias propagation.
+5 Datasets
+We evaluate our framework on three benchmark networks of varying size and realism.
+Synthetic Adult-Based Network.
+We construct a semi-synthetic dataset based on the UCI Adult dataset (Kohavi et al.
+1996), a standard benchmark in fairness research.
+The graph includes 15 original variables, with causal edges informed by prior work (Kilbertus et al.
 2017; Nabi and Shpitser 2018).
-The DAG assumes that demographic features (e.g., race, sex) influence education and occupation; education affects income both directly and via mediators like occupation and capital gains; and age and marital status serve as confounders or mediators in employment-related paths.
-Capital-related features are modeled as downstream effects of occupational history, contributing to income.
-To simulate structural unfairness, we inject direct edges from sex and race to income, consistent with causal fairness benchmarks (Nabi and Shpitser 2018; Kilbertus et al.
-2017).
-These pathways are complemented by indirect paths (e.g., sex $\rightarrow$ education $\rightarrow$ occupation $\rightarrow$ income) to facilitate the estimation of direct, indirect, and total effects in fairness evaluation.
-The baseline graph and technical details on the data generation process can be found in Appendix A.1.
-5.1 Structured Noise Augmentation
-To test robustness under realistic conditions, we augment the clean synthetic data with structured noise and confounding, emulating sources of measurement error and label corruption common in real-world datasets.
-Additive Noise for Continuous Variables.
-Gaussian noise is added to all continuous variables post-generation to simulate measurement variability.
-For each continuous variable X, noise is sampled as:
-$X_{i}^{\text{noisy}}=X_{i}+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{X}^{2}),\quad\sigma_{X}=\alpha\cdot\mathrm{std}(X)$ (8)
-where $X_{i}$ is the original value for sample $i$, $\epsilon_{i}$ is a random error drawn from a normal distribution with mean 0 and standard deviation $\sigma_{X}$, $\sigma_{X}=\alpha\cdot\mathrm{std}(X)$, where $\alpha\in[0,1]$ is a tunable noise scaling factor set to $\alpha=0.1$.
-This maintains proportionality between noise magnitude and feature variability, consistent with additive noise models used in ICA-LiNGAM and structural equation modeling (SEM) frameworks (Shimizu et al.
-2006; Bollen 1989).
-Symmetric Label Noise in Categorical Variables
-To emulate mislabeling, we introduce symmetric label noise into categorical variables.
-For each categorical variable, a proportion $\gamma$ of values (5%) are randomly replaced with an alternative category, drawn uniformly from the remaining labels.
-For each categorical variable $Y$ with $K$ classes (e.g., ”Male”, ”Female”), uniform symmetric noise is introduced using the following equation.
-Let $\gamma\in[0,1]$ be the label noise probability. Then for each sample $i$:
-$P(\tilde{Y}_{i}=y^{\prime})=\begin{cases}1-\gamma&\text{if }y^{\prime}=Y_{i}\\ \frac{\gamma}{K-1}&\text{if }y^{\prime}\neq Y_{i}\end{cases}$ (9)
-Where $Y_{i}$ is the original category label, $\tilde{Y}_{i}$ is the observed (possibly noisy) label, $K$ is the number of possible categories for $Y$, and $y^{\prime}\in\mathcal{Y}$ are all possible class values.
-This means with probability $\gamma$, the label is flipped uniformly to one of the $K-1$ incorrect categories.
-And with probability $1-\gamma$, the label remains correct.
-This uniform corruption model known as symmetric label noise, is widely used in label-noise robust learning (Natarajan et al.
-2013; Patrini et al.
-2017).
-Latent Confounding via Unobserved Variable $U$
-To simulate unmeasured social influences (e.g., privilege or institutional access), we introduce a latent confounder $U\sim\mathcal{N}(0,1)$ that affects intermediate variables such as years of education, capital gain, and work hours, thereby indirectly influencing income.
-This reflects latent societal advantages that are not captured in observed features.
-The latent confounder is injected via additive linear terms into structural equations following standard assumptions in structural causal models (SCMs).
-For the binary outcome income, we use a nonlinear threshold function that combines observable and latent contributions.
-This setup reflects bias amplification near decision boundaries and is consistent with prior fairness simulation frameworks.
-Full equations and implementation details are provided in Appendix A.2.
-Controlled Replication Across Seeds
-To evaluate generalization and fairness robustness, we generate four datasets using different random seeds, each with the same underlying graph but distinct noise realizations.
-Noise is added post-graph generation, preserving causal structure while introducing realistic variability.
-This mirrors standard practices in fairness benchmarking and robustness testing for CD.
-6 Datasets
-To evaluate the performance of the proposed LLM-based CD framework with AL and dynamic scoring, a series of experiments are conducted on benchmark networks.
-These networks are described in the following.
+Demographics (e.g., race, sex) influence education and occupation, which in turn affect income, both directly and via mediators such as capital gains.
+Age and marital status act as confounders or mediators.
+To simulate structural bias, we inject direct edges from race and sex to income, alongside indirect paths (e.g., sex $\rightarrow$ education $\rightarrow$ income).
+These allow evaluation of direct, indirect, and total effects.
+Details on graph construction and data generation are in the Appendix.
 Child Causal Network.
-The Child network is a medium-scale Bayesian network simulating causal relationships relevant to congenital heart disease in newborns (Spiegelhalter et al.
+A 20-node, 25-edge Bayesian network modeling clinical, environmental, and parental factors in congenital heart disease (Spiegelhalter et al.
 1993).
-It includes 20 nodes and 25 edges spanning medical, parental, and environmental factors.
-Variables include both categorical and continuous types, covering clinical features such as birth asphyxia, lung function, cardiac mix, chest X-ray results, and outcomes like mortality and hospital stay.
-Neuropathic Pain Causal Network.
-The Neuropathic Pain Diagnosis Simulator (Tu et al.
-2019) models causal relationships underlying neuropathic pain using a graph of 221 nodes and 770 edges, based on real patient data.
-It includes symptom, pattern, and pathophysiological diagnoses from actual patient data.
-As these benchmarks lacks sensitive variables, we do not assess fairness; evaluation focuses solely on structural accuracy.
+Variables include both categorical and continuous types.
+Neuropathic Pain Network.
+A large-scale clinical graph with 221 nodes and 770 edges capturing pathophysiological and symptomatic relationships in neuropathic pain diagnoses, derived from real-world patient data (Tu et al.
+2019).
+Since the latter two datasets lack sensitive attributes, we focus on structural accuracy rather than fairness.
 6 Experiments
-To evaluate the proposed LLM-based CD framework, we conduct a series of experiments on the introduced datasets.
-The primary goal is to assess the framework’s ability to improve causal graph accuracy and recover fairness-relevant pathways under varying configurations and noise conditions.
-This section outlines our experimental design, evaluation metrics, and baseline comparisons.
-7.1 Baseline Methods
-We compare our method against the following CD approaches:
-Peter-Clark (PC) Algorithm.
-A constraint-based method using conditional independence tests to iteratively prune edges from a fully connected graph, followed by orienting edges to form a DAG.
-It assumes causal sufficiency and the faithfulness assumption, relying heavily on accurate independence tests (Spirtes and Glymour 1991).
-Greedy Equivalence Search (GES) Algorithm.
-A score-based method that operates in two phases: a forward phase that incrementally adds edges to improve a score function (e.g., BIC), and a backward phase that removes edges to optimize the graph structure further.
-It assumes causal sufficiency and focuses on equivalence classes of DAGs to improve computational efficiency (Meek 1997).
-NOTEARS Algorithm.
-Formulates CD as a continuous optimization problem by introducing a differentiable acyclicity constraint.
-This allows it to learn DAGs using gradient-based optimization, balancing data fit and sparsity (Zheng et al.
-2018).
-DAGMA Algorithm.
-Integrates neural networks and optimization techniques to model nonlinear causal relationships.
-It uses a masked autoencoder architecture to enforce sparsity and incorporates a differentiable acyclicity condition to ensure that the resulting graph is a valid DAG (Bello, Aragam, and Ravikumar 2022).
-LLM-based Pairwise Method.
-Utilizes LLMs to infer causal relationships through pairwise queries based on metadata and variable descriptions. While intuitive, it scales poorly with graph size (Kıcıman et al. 2023).
-LLM-based BFS Method.
-LLM-based BFS appraoch introduced by Jiralerspong et al. (Jiralerspong et al. 2024) which was discussed in the Sections 1 and 3.
-All baselines are implemented using the open-source framework from (Jiralerspong et al. 2024) for consistency.
-7.2 Experimental Design
-We apply Bayesian optimization with a Gaussian Process (GP) surrogate model to explore the influence of key hyperparameters on performance.
-This enables tuning of our query scoring and LLM prompting strategy to enhance edge discovery and fairness-related signal recovery (Mockus 1994; Snoek, Larochelle, and Adams 2012).
-Optimization is performed using gp_minimize from scikit-optimize.
-The hyperparameters optimized include:
-  • Scoring Weights: Relative contributions of MI, PCorr, and query history in Equation 1.
-  • Score threshold: Minimum dynamic score for querying a variable pair.
-  • LLM Temperature: Controls the stochasticity of LLM responses.
-  • Max Iterations: Maximum number of AL iterations per trial.
-To evaluate robustness, we divide trials into chunks that sample different regions of the hyperparameter space.
-Each chunk records graph outputs and performance metrics, and the top-performing configuration from each run is used for final reporting.
-Full details on the hyperparameters used in each set of experiments are provided in Appendix B.
-We evaluate all methods on four synthetic datasets generated from the Adult-based causal graph (Section 5) using different random seeds, which preserve the underlying structure while introducing independent data realizations.
-Performance metrics are averaged across these seeds.
-Fairness pathway analyses are conducted only on the Adult-based datasets, focusing on paths from sensitive attributes (sex, race, age) to the outcome variable (income).
-We also evaluate all methods on the Neuropathic and Child benchmark networks, averaging results for LLM-based methods across five independent runs.
-7.3 Performance Evaluation Metrics
-We assess graph reconstruction performance using standard structure-learning metrics measured against the ground-truth DAG.
-Precision and Recall.
-$\text{Precision}=\frac{\text{TP}}{\text{TP + FP}},\quad\text{Recall}=\frac{\text{TP}}{\text{TP + FN}}$
-F1 Score.
-[…]
-Edge Count Comparison.
-We compare the number of predicted edges with the number in the ground-truth graph.
-DAG Validation.
-Checks acyclicity of the predicted structure.
-Normalized Hamming Distance (NHD).
-$\text{NHD}=\frac{\text{Mismatches}}{n^{2}}$
-where n refers to the number of nodes in the graph.
-Adjacency Accuracy.
-Accuracy=Correct $\text{Accuracy}=\frac{\text{Correct entries}}{n^{2}}$
+We evaluate our LLM-based CD framework across the benchmark datasets, focusing on graph accuracy and fairness-relevant path recovery under varying noise and hyperparameter settings.
+6.1 Baselines
+We compare against the following methods:
+  • PC Algorithm (Spirtes and Glymour 1991): Constraint-based method using independence tests.
+  • GES (Meek 1997): Score-based approach optimizing a score function (e.g., BIC) via forward-backward search.
+  • NOTEARS (Zheng et al. 2018): Continuous optimization using a differentiable acyclicity constraint.
+  • DAGMA (Bello, Aragam, and Ravikumar 2022): Neural network-based CD with sparsity and acyclicity constraints.
+  • LLM Pairwise (Kıcıman et al. 2023): Uses LLM to infer pairwise causality from metadata.
+  • LLM BFS (Jiralerspong et al. 2024): Explores graph structure via LLM-guided breadth-first querying.
+All baselines are run using the open-source code from (Jiralerspong et al. 2024).
+6.2 Experimental Setup
+We use Bayesian optimization (gp_minimize) with a Gaussian Process surrogate to tune key hyperparameters (Mockus 1994; Snoek, Larochelle, and Adams 2012):
+  • Scoring Weights: Weights for MI, PCorr, and query history (Eq. 1)
+  • Score Threshold: Minimum score required for querying a pair
+  • LLM Temperature: Controls randomness in responses
+  • Max Iterations: AL rounds per run
+Each query uses a multi-turn GPT-4 (0125-preview) prompt with prior discoveries and variable metadata (Section 3).
+A directed edge $X\rightarrow Y$ is added if the reply contains <Answer>Yes</Answer> and does not create a cycle.
+If ambiguous, no edge is added.
+Confidence scores are derived from the average probability of top-5 tokens based on log-probs; defaulting to 0.5 if unavailable.
+Trials are divided into chunks across hyperparameter space. We report results from the best-performing configuration for each run. Full prompt design and search ranges are detailed in the Appendix.
+We evaluate all methods on four synthetic variants of the Adult-based dataset (Section 5), varying random seeds to preserve the underlying structure while introducing independent data realizations.
+Fairness analysis targets paths from sex, race, and age to income.
+Structural accuracy is also assessed on the Child and Neuropathic networks, with LLM-based methods averaged over five independent runs.
+6.3 Evaluation Metrics
+We assess reconstruction accuracy using standard metrics:
+  • Precision & Recall: Defined as $\text{Precision}=\frac{\text{TP}}{\text{TP + FP}}$ and $\text{Recall}=\frac{\text{TP}}{\text{TP + FN}}$.
+  • F1 Score: Harmonic mean of precision and recall.
+  • Edge Count: Number of predicted edges vs. ground truth.
+  • Acyclicity: Ensures output is a valid DAG.
+  • Normalized Hamming Distance (NHD): $\text{NHD}=\frac{\text{Mismatches}}{n^{2}}$, where $n$ is the number of nodes.
+  • Adjacency Accuracy: $\text{Accuracy}=\frac{\text{Correct entries}}{n^{2}}$.
 7 Results and Analyses
-8.1 Causal Discovery Performance Comparison Between Methods
-Table 1: Performances of Methods on the Adult-based Synthetic, Child and Neuropathic Causal Networks
+7.1 Graph Performance Comparison
 Adult-based Synthetic Method Acc. ($\uparrow$) F1 Score ($\uparrow$) Precision Recall NHD ($\downarrow$) NHD Ratio ($\downarrow$) # Pred. Edges
@@
 Proposed Method 0.073 0.136 0.690 0.075 0.109 0.864 84
-Table 1 summarizes performance across the synthetic Adult-based, Child, and Neuropathic networks using accuracy, precision, recall, F1 score, and NHD.
-Across all datasets, the proposed method consistently outperforms baseline approaches in F1 score, precision, and structural alignment with the ground-truth graph.
-The proposed method generally leads to conservative graph reconstruction compared to the baselines.
+Table 1: Performances of Methods on the Adult-based Synthetic, Child, and Neuropathic Causal Networks.
+All reported values are averaged over four random seeds.
+Metrics exhibit minimal variance across runs ($\pm 2\times 10^{-4}$ in F1 score).
 Synthetic Adult-based Network.
-The proposed method outperforms all baselines in most metrics except recall.
-It achieves the highest F1 score (0.585), accuracy (0.413), precision (0.792), and the lowest NHD (0.109) and NHD ratio (0.415).
-This demonstrates its superiority in structural accuracy and edge prediction while producing a more concise graph (17 predicted edges) than other methods.
-LLM Pairwise method achieves the highest recall (0.813) but overpredicted edges (69).
-Classical methods like GES and PC achieved moderate F1 scores (0.473 and 0.382, respectively), but their graphs had higher NHDs and more structural violations.
-Continuous optimization methods (NOTEARS and DAGMA) underperform across most metrics, indicating sensitivity to noise and difficulty capturing fairness-relevant structure.
-Child Network (20 nodes, 25 edges).
-The proposed method achieves the highest F1 score (0.533), precision (0.601), and accuracy (0.364), outperforming both statistical and LLM-based baselines.
-It also achieves the lowest NHD ratio (0.467), indicating better alignment with the ground-truth structure.
-Although NOTEARS achieves a slightly better raw NHD, our method offers a better precision-recall balance.
-The LLM Pairwise method shows strong recall (0.559) but overpredicts edges, leading to lower precision and weaker structural alignment.
-Neuropathic Network (221 nodes, 770 edges).
-The proposed method achieves the highest F1 score (0.136), the highest precision (0.690), and the lowest normalized Hamming distance (NHD) ratio (0.864) across all tested methods.
-While these metrics are relatively low in absolute terms, they reflect a broader trend where all evaluated methods struggle to recover the full causal structure in this large, high-dimensional graph.
-This underscores the inherent difficulty of CD in settings characterized by sparse, noisy dependencies and complex variable interactions.
-Still, the proposed method exhibits higher performance compared to baseline approaches by achieving a better balance between edge prediction accuracy and false positive control.
-The LLM-BFS baseline fails entirely on this dataset (F1 = 0, accuracy = 0), likely due to its static query strategy being overwhelmed by the Neuropathic graph’s high dimensionality and sparsity.
-This underscores the limitations of fixed traversal methods in large-scale settings.
-In contrast, our method’s use of dynamic scoring and AL improves query prioritization, enabling more effective discovery under complex conditions.
+Our method achieves top scores across F1 (0.585), accuracy (0.413), precision (0.792), and NHD (0.109), recovering a compact graph (17 edges) with strong alignment to ground truth.
+LLM Pairwise attains high recall (0.813) but overpredicts, while GES and PC show moderate F1 but higher NHDs.
+Optimization-based methods struggle with fairness structure.
+Child Network.
+Again, our method leads in F1 (0.533), accuracy (0.364), and NHD ratio (0.467), balancing edge precision and coverage.
+NOTEARS achieves slightly better raw NHD but worse recall.
+LLM Pairwise finds many indirect effects, but at the cost of low precision.
+Neuropathic Network.
+Despite the challenge of 221 nodes, our method yields highest F1 (0.136) and precision (0.690), outperforming baselines.
+All baselines degrade sharply; LLM BFS fails entirely (F1 = 0).
+Dynamic scoring enables better query selection under sparsity.
 Scalability.
-In addition to performance, we also note qualitatively that our method scales more efficiently than optimization-based approaches (e.g., DAGMA, NOTEARS), and the LLM-pairwise method while offering greater structural fidelity.
-We also perform a complexity analysis described in Appendix E showing how our proposed method could theoretically scale better than the LLM BFS method on significantly large sparse networks, characteristics which real-world networks usually exhibit.
-This balance that the proposed method provides is evident in the Neuropathic setting, where computationally intensive baselines suffer performance degradation or scalability bottlenecks.
-Due to these scalability and resource limitations, experiments using the GES and LLM pairwise methods were not completed on the Neuropathic graph.
-Reproducibility and Implementation Discrepancies.
-We observe lower performance for LLM-BFS and LLM-Pairwise compared to results reported in Jiralerspong et al.
-(2024), particularly on the Child and Neuropathic networks.
-This may be due to (1) non-determinism and ongoing updates to models like GPT-4, which may impact consistency across evaluations, especially in semantically ambiguous cases.
-(2) implementation differences in token limits or message handling, (4) lack of access to original variable descriptions for the Neuropathic dataset, affecting metadata quality.
-Despite these differences, our implementation adheres to the BFS protocol and reveals its limitations in high-dimensional settings, further motivating our enhancements through dynamic scoring and AL.
-8.2 Fairness Pathway Recovery and Bias Attribution
-We assess each method’s fairness-awareness by analyzing causal paths from sex, age, and race to income in the Adult-based synthetic network.
-The ground-truth graph contains 2 direct and 25 indirect fairness paths, with most influence mediated (Indirect = 0.93, Direct = 0.07).
-A total effect (TE) of 4.89 and normalized bias (C_bias = 28.46) indicate substantial structural bias, aligning with real-world disparities (Appendix D).
-We compare each method’s recovered graph in terms of these fairness-aware path metrics.
-A table detailing the averages of all the fairness and path-related scores for the methods across seed runs is included in Appendix D, as well as a visualization of the graph generated by the outputs of the proposed method.
-Figure 2 shows the number of direct, indirect, and spurious paths per method.
-Due to the wide variance where some methods generate fewer than 5 paths (e.g., NOTEARS), others over 600 (e.g., GES), we use a log-scaled y-axis to preserve readability and comparability without distorting proportions.
-Figure 2: Log-scaled path type counts (direct, indirect, spurious) across methods. The log scale enables fair comparison by compressing the large variation in total path counts.
-Our proposed method is the only approach to correctly identify both direct fairness paths from sex and race to income, while correctly omitting a spurious direct path from age, demonstrating both sensitivity to true bias and restraint against overgeneralization.
-In contrast, LLM Pairwise and BFS methods introduce an incorrect direct path from age, yielding three direct paths in total.
-This inflated their $C_{\text{bias}}$ values to appear closer to the baseline but was driven by over-attribution rather than accurate recovery.
-These methods tend to over-attribute causal influence to demographically meaningful variables due to LLM priors and lack of structural constraints, which can amplify perceived bias.
-Classical statistical methods, including PC, GES, NOTEARS, and DAGMA, largely underperform.
-DAGMA and NOTEARS fail to identify any valid paths (direct or indirect) from the sensitive attributes to income, producing exclusively spurious paths.
-As a result, their $C_{\text{bias}}$ values are near zero, reflecting an inability to detect embedded social bias mechanisms, likely due to their strict dependence on data-driven associations and conservative edge inclusion thresholds.
-PC identifies few fairness-relevant paths, and missed both direct edges, leading to a low bias score.
-This aligns with its constraint-based nature, which can be overly sensitive to noise in independence tests, particularly with sparse or weak signals.
-GES, while slightly better at recovering indirect effects, also fails to identify direct paths.
-However, due to its aggressive edge addition during the forward search phase, it captured a larger number of fairness-related pathways overall, majority of which are spurious.
-This explains its relatively close $C_{\text{bias}}$ score to the baseline, though the accuracy of its causal assignments remains debatable.
-The graphs constructed by all methods are shown in Appendix D.
-Overall, our proposed approach demonstrates a comparatively strong alignment with the fairness-relevant structure of the ground truth.
-It recovered both true direct effects from sex and race to income, while avoiding spurious paths such as a direct effect from age, which is introduced erroneously by other methods.
-Although the resulting $C_{\text{bias}}$ value is lower than the baseline, this may reflect the method’s prioritization of high-confidence, high-scoring queries and its dynamic scoring mechanism that discourages over-inclusion.
-This more conservative behavior potentially reduces the risk of false positives in fairness-sensitive settings, though it may also underestimate total bias when indirect effects are more diffuse.
-In addition, our method also exhibits the most stable performance across runs, as reflected by its consistently low standard deviation on key metrics such as indirect path count, total effect, and normalized bias as seen in the table in Appendix D.
-This reliability is important in fairness-sensitive contexts, where reproducibility and robustness are critical.
-Figure 3: Comparison of normalized bias ($C_{\text{bias}}$) across methods.
-8.3 Parameter Impact Analysis on Proposed Method
-We analyze the Bayesian optimization trials using a Random Forest regressor to identify which hyperparameters most influence F1 performance.
-Max iterations have the greatest impact, followed by score threshold and temperature (Figure 4).
-This suggests that deeper exploration in AL improves graph construction by reducing false negatives.
-Figure 5 supports this, showing F1 scores increasing with iteration budget, up to a plateau, emphasizing the trade-off between computational cost and discovery depth.
-Figure 4: Feature importance of hyperparameters in predicting F1 score, computed via Random Forest regression. max_iterations has the highest predictive influence.
-Figure 5: Relationship between max_iterations and F1 score. Higher iteration budgets generally yield better causal graph construction performance.
-The correlation matrix (Figure 6) shows strong negative correlations among scoring weights, reflecting their competitive role in the dynamic scoring mechanism, while only max iterations correlates positively with F1.
-This suggests that, for smaller networks like the Adult-based synthetic graph, exploration depth outweighs weight tuning in driving discovery performance.
-Figure 6: Pearson correlation heatmap between F1 score and key hyperparameters. Notably, max iterations shows strong positive correlation with F1, while weight terms are moderately anti-correlated.
-As graph size increases, parameter importance shifts.
-In the large Neuropathic network (221 nodes), statistical weights show the strongest positive correlation with F1 (Appendix C), suggesting that MI and PCorr are crucial for narrowing the query space in high-dimensional settings.
-Conversely, LLM confidence and query history weights become less influential, likely due to semantic cues being less reliable across deeper or denser graphs.
+Our method scales better than NOTEARS, DAGMA, or LLM Pairwise while offering greater structural fidelity.
+Complexity analysis (detailed in Appendix) shows favorable scaling theoretically in large sparse graphs, common in real-world networks.
+GES and LLM Pairwise were excluded from Neuropathic due to resource limits.
+Reproducibility.
+LLM-BFS and Pairwise underperform vs.
+Jiralerspong et al.
+(2024), likely due to (1) GPT-4 updates and nondeterminism, (2) token limits or message handling, and (3) lack of access to original variable descriptions for the Neuropathic dataset, affecting metadata quality.
+Our implementation adheres to the BFS protocol and exposes its limitations under scale, motivating our enhancements.
+7.2 Fairness Pathway Recovery
+We evaluate fairness path recovery on the Adult-based synthetic graph, where the true structure contains 2 direct and 25 indirect paths from sex, age, and race to income.
+TE = 4.89 and $C_{\text{bias}}=28.46$.
+Figure 2: Log-scaled path counts (direct, indirect, spurious) per method.
+Our method uniquely recovers both true direct paths (from sex and race) while omitting the spurious age path, showing strong fairness alignment.
+LLM methods over-attribute (3 direct paths) due to LLM priors, inflating $C_{\text{bias}}$.
+PC and GES miss direct effects; NOTEARS and DAGMA capture none, yielding near-zero $C_{\text{bias}}$.
+Figure 3: Normalized bias ($C_{\text{bias}}$) across methods.
+Overall, our method prioritizes high-confidence discovery, avoiding false positives and maintaining robustness across seeds as reflected by its consistently low standard deviation on key metrics such as indirect path count, total effect, and normalized bias as seen in the table in Appendix.
+Its conservative nature may underestimate bias but enhances reliability in fairness-critical settings.
+7.3 Ablation and Parameter Influence
+Bayesian optimization trials were analyzed via Random Forest regression to assess parameter impact on F1 score.
+Figure 4: Hyperparameter importance on F1 score.
+Findings.
+Max iterations have the strongest effect, suggesting deeper AL improves recovery.
+Score threshold and temperature also influence performance.
+Correlation analysis (Figure 5) shows anti-correlation among scoring weights, indicating competitive trade-offs.
+In small graphs, deeper querying matters more; in large ones (e.g., Neuropathic), MI/PCorr dominate (details in Appendix).
 This highlights how our dynamic scoring mechanism could adapt to scale, favoring semantic guidance in small graphs and statistical grounding in larger ones.
+Figure 5: F1 vs. hyperparameter correlation (Adult).
 8 Discussion
-Our findings show that augmenting LLM-based CD with AL and dynamic scoring improves both structural accuracy and fairness sensitivity.
-The proposed method consistently outperformed baselines across the Adult-based synthetic network, the Child graph, and the large-scale Neuropathic Pain network.
-The dynamic scoring mechanism enables effective trade-offs between statistical dependencies and LLM-derived semantics, adapting to graph complexity.
-This is especially valuable in high-dimensional settings, where statistical cues helped narrow the query space.
-AL further optimized discovery by prioritizing informative, non-redundant queries.
-In terms of fairness, our method is the only one to recover both direct paths from sensitive attributes (sex, race) to income while avoiding false positives (e.g., a spurious link from age).
-This indicates strong potential for fairness audits with greater precision and reduced over-attribution.
-Parameter sensitivity analysis reveals that exploration budget (e.g., max iterations) is the dominant driver of performance.
-While LLM signals aid early search stages, statistical cues grow in importance with graph size, highlighting the adaptive nature of our hybrid scoring framework across discovery contexts.
-While our findings highlight the general potential of CD methods to inform fairness assessments, they also underscore a critical caveat where such methods, particularly when applied naively, can introduce spurious pathways or erroneous connections involving sensitive attributes.
-Across baseline methods, we observe frequent over or under attribution of causal influence to variables like race, sex, or age, leading to inflated or misleading estimates of structural bias.
-This is especially problematic in fairness contexts, where false positives (e.g., wrongly inferring a direct path from a sensitive attribute to an outcome), as indicated by the misleadingly similar C_bias scores of graphs from non-performing CD methods to the baseline, may trigger unjustified policy interventions or obscure true sources of inequity.
-The high variance in graph structure suggests that CD outputs should be interpreted with considerable caution.
-Practitioners must validate inferred relationships using domain knowledge, robustness checks, and sensitivity analyses before drawing conclusions about discrimination or bias.
-Without such safeguards, CD risks becoming a source of misinformation rather than a tool for mitigating bias.
+Our findings demonstrate that combining LLM-based CD with AL and dynamic scoring improves both structural accuracy and fairness sensitivity across diverse datasets.
+The proposed method consistently outperforms statistical, LLM-only, and hybrid baselines on both small and large networks.
+Dynamic scoring balances LLM semantic priors with empirical signals, adapting to graph complexity and data quality.
+In early iterations, LLM judgments guide exploration when statistical signals are weak.
+As more information is gathered, mutual information and partial correlation become dominant, reflecting a shift from semantic exploration to empirical refinement.
+AL enhances efficiency by prioritizing uncertain, low-redundancy variable pairs, avoiding the limitations of fixed-order strategies.
+Our method uniquely recovers fairness-critical paths, e.g., sex $\rightarrow$ education $\rightarrow$ income—while avoiding spurious ones like age $\rightarrow$ income, often introduced by LLM-only baselines.
+This improves fairness diagnostics for impact assessments in domains like hiring or lending.
+While our fairness evaluation relies on a semi-synthetic benchmark, this is necessary as real-world datasets lack ground-truth causal graphs for sensitive attributes and outcomes.
+However, our results also underscore broader concerns: CD methods, including ours, may introduce spurious or incomplete structures if used without scrutiny.
+In fairness contexts, such errors carry high stakes.
+False positives (e.g., unjustified bias paths) can trigger overcorrections, while false negatives obscure actual harms.
+We find that even underperforming methods can appear fairness-aware using naive metrics like $C_{\text{bias}}$, which may be inflated by erroneous causal paths.
+This highlights the need for path-level interpretability rather than summary statistics alone.
+To mitigate these risks, our approach integrates LLMs as prior-informed agents within a selective AL loop.
+This reduces unnecessary queries, avoids early commitment errors, and improves efficiency over fixed-query strategies such as BFS.
+Like traditional DAG methods, our framework embraces uncertainty and treats CD as hypothesis generation.
+While it is argued that LLMs may not perform causal reasoning per se, we treat them as semantic priors filtered through statistical signals and query history.
+This is analogous to expert elicitation—subject to bias, but useful when paired with empirical validation.
+As argued by (Imbens 2020) and (McCoy et al.
+2023), epistemic uncertainty is intrinsic to CD.
+LLM-guided proposals can still surface plausible structures, especially in fairness-critical settings.
+Prior work has emphasized this role (Kıcıman et al.
+2023); we build on it with a dynamic querying process that improves recovery while reducing false positives in bias pathways.
+Our broader aim is to support interpretable, efficient, and socially aware fairness diagnostics, offering complementary insight to notions like demographic parity or counterfactual fairness.
+By uncovering how sensitive variables influence outcomes through direct and mediated paths, our method moves beyond outcome disparities toward causal insight.
+This aligns with calls for path-level transparency in AI audits from fairness, accountability, and legal communities.
+We see this work as a step toward embedding causal reasoning into policy evaluations, model and risk assessments, tools needed to build more trustworthy socio-technical systems.
+This approach provides actionable insights for non-technical stakeholders.
+By surfacing fairness-relevant pathways, the method enables practitioners, compliance teams, or policy analysts to trace how sensitive attributes indirectly shape outcomes.
+These insights can guide interventions, from revising decision rules to informing accountability processes in hiring, lending, or regulatory oversight.
 9 Limitations
-Simplified Causal Modeling in Synthetic Data.
-Our synthetic graph, based on the UCI Adult dataset, models income as a terminal node to isolate fairness-related pathways.
-This design aids in evaluating methods’ ability to detect direct and indirect influences on income.
-However, it omits downstream effects that income may have in reality—such as on health, wealth, or opportunity, thus excluding feedback dynamics and longitudinal consequences.
-Future work could extend this framework to capture post-outcome influences and assess fairness in more dynamic causal settings.
-Hyperparameter Sensitivity and Computational Constraints.
-Although effective, our method requires careful tuning of multiple hyperparameters (e.g., scoring weights, thresholds, LLM temperature), demanding domain expertise.
-LLM-based querying also incurs substantial computational cost, especially for large networks.
+Simplified Synthetic Causal Structure.
+Our synthetic graph models income as a terminal node to isolate fairness pathways, aiding evaluative clarity.
+However, this excludes realistic downstream effects (e.g., on health, opportunity), omitting feedback or temporal dynamics.
+Future work should extend this to longitudinal or post-outcome settings.
+Hyperparameter Sensitivity and Compute Costs.
+Our method requires tuning multiple hyperparameters (e.g., score weights, temperature), which can be resource-intensive.
+LLM querying is computationally costly, particularly for large graphs like Neuropathic.
 Token limits restrict the ability to encode full metadata, particularly in complex graphs like Neuropathic, leading to incomplete or inconsistent responses.
 Specialized variable names may be misinterpreted, especially at higher temperatures, introducing semantic errors and prediction noise.
-These factors collectively pose scalability and usability challenges, particularly for resource-constrained settings or high-dimensional domains.
-Reproducibility Challenges in LLM-Based Causal Discovery.
-Discrepancies between our results and those reported by Jiralerspong et al.
-(2024) highlight broader reproducibility concerns in LLM-based CD.
-LLMs are inherently stochastic and sensitive to prompt design, versioning, query order, and API behavior, factors rarely standardized or reported.
-Additionally, hidden dependencies on tokenization, temperature, or latent state carryover can affect outcomes, making benchmarking difficult.
-These challenges underscore the need for greater transparency, reproducible prompt templates, and controlled evaluation protocols in future work.
-Risk of Social Bias in LLM Responses.
+These factors collectively pose scalability and usability challenges, particularly for resource-constrained settings.
+Reproducibility of LLM-based CD.
+While our method is LLM-agnostic, we observe discrepancies with results from Jiralerspong et al.
+(2024), highlighting broader reproducibility challenges.
+LLMs are stochastic and sensitive to prompt design, versioning, query order, and API behavior, factors rarely standardized or reported.
+Without standardized prompting protocols or access to latent states, consistent benchmarking remains difficult.
+Risk of Social Bias in LLMs.
 LLMs trained on web-scale data may encode social biases, leading to causal inferences that reflect stereotypes rather than ground-truth mechanisms.
 This risks semantic hallucinations and biased edge assignments, particularly in fairness-critical applications.
-Domain Dependency and Metadata Limitations.
-A further limitation lies in the domain dependency of our approach.
-Because our framework relies on variable metadata to guide LLM-based querying, it is best suited to domains with well-defined, semantically meaningful feature names (e.g., healthcare, census, education).
-In settings where metadata is sparse, ambiguous, or domain-specific (e.g., sensor networks or high-dimensional genomics), LLM-based judgments may degrade or hallucinate structure (E.g.
-Neuropathic network).
+This risk is addressed through statistical weighting and confidence-based filtering, as described in Section 3.
+Domain and Metadata Dependence.
+Our framework depends on interpretable variable metadata to guide LLM queries.
+It performs best in domains like healthcare or census data.
+In sparse, technical, or ambiguous domains (e.g., genomics, sensor data), performance degrades.
+While LLM priors scale efficiently, human-in-the-loop priors could be incorporated for low-interpretability domains.
 Future work could explore hybrid prompting strategies or domain adaptation techniques to extend semantic CD to less interpretable or evolving data environments.
 10 Conclusion
-In this work, we propose a fairness-driven CD framework that integrates LLMs with AL and dynamic scoring.
-Our method outperforms baselines in recovering accurate causal structures across various benchmark networks, and demonstrates more conservative behavior in identifying fairness-relevant pathways, reducing false positives in bias attribution.
-By combining statistical signals with semantic priors, and prioritizing high-informative queries, the proposed framework advances the reliability and fairness-awareness of LLM-based CD.
-To support robust evaluation, we introduce a semi-synthetic benchmark based on the UCI Adult dataset, embedding fairness-relevant causal structure, confounding, and noise.
-This benchmark provides a controlled yet realistic testbed for future fairness-oriented CD research.
-Our hyperparameter analysis revealed that exploration depth (max iterations) strongly influences performance, while the importance of statistical versus LLM-derived signals varies with graph scale.
-While our approach improves recovery of fairness-relevant pathways, our framework does not directly translate into end-to-end fairness guarantees in downstream tasks.
-Bridging the gap between discovery and actionable fairness interventions remains an important area for future integration.
-Other future directions include expanding to temporal graphs, refining prompt engineering for domain-specific metadata, and building reproducible pipelines for benchmarking LLM-guided causal inference.
+We propose a fairness-driven CD framework that integrates LLMs with AL.
+Across benchmarks including synthetic and semi-synthetic real-world networks, our method outperforms prior baselines in structural accuracy and reduces false positives in fairness-critical path recovery.
+By combining statistical dependencies with LLM-based semantic priors and prioritizing informative queries, our framework improves both reliability and fairness-awareness in CD.
+To support reproducible evaluation, we introduce a semi-synthetic benchmark based on the UCI Adult dataset, embedding structural bias, latent confounding, and noise.
+This provides a realistic yet controlled testbed with ground-truth graphs for consistent comparison across CD methods.
+Our analysis shows that exploration budget (e.g., number of iterations) significantly affects performance, with the relative importance of LLM-derived versus statistical signals shifting with graph size and complexity.
+While our method improves fairness-path recovery, it does not offer end-to-end guarantees for downstream fairness outcomes.
+Instead, it uncovers explainable causal pathways, particularly those involving sensitive attributes intended to guide audits, impact assessments, or stakeholder review.
+This positions our framework as an improvement in diagnostic tools for fairness evaluations.
+Use cases include hiring audits, policy design, and regulatory compliance reporting.
+By recovering interpretable causal paths, our method offers a step toward tools that support fairness-focused exploration, potentially informing future integration into dashboards or algorithmic auditing pipelines.
+We plan to extend the framework to dynamic graphs and release open-source pipelines for scalable, reproducible fairness audits.
 Ethical Statement.
@@
 Future work should explore more efficient, bias-aware models to reduce this impact.
-Appendix A Appendix: Synthetic Data Generation Rules
-This section provides more detail on the Adult-based synthetic data general rules, formulas, and process.
+Acknowledgments
+This study was funded by NSF #2047296 and OpenAI’s Researcher Access Program.
+Appendix A Synthetic Data Generation
+This section provides more detail on the Adult-based synthetic data general rules, formulas, and process.The graph was constructed based on causal assumptions documented in prior fairness research (Kilbertus et al.
+2017; Nabi and Shpitser 2018).
+Variable relationships were chosen to reflect widely accepted socioeconomic patterns and to ensure transparency and reproducibility, minimizing author-imposed bias.
 A.1 Data Generation Process
@@
 relationship Categorical (4 classes) marital status 0 if marital status==1; 1 if marital status==3; 2 if marital status==0; else 3
-[…]
-[…]
-[…]
-[…]
+capital gain Numerical occupation $\mathcal{N}(5000+1500\cdot U,\ 2000)$ if occupation==1; else 1000 + 500 $\cdot$ U or 0
+capital loss Numerical occupation $\mathcal{N}(1000,\ 500)$ if occupation==1; else 300 or 0
+work hours per week Numerical occupation, sex $\mathcal{N}(45+2\cdot U,\ 5)$ if sex==0; else 40 + U or 35
+final weight Numerical None $\mathcal{N}(180000,\ 20000)$
 income Categorical (¡=50K, ¿50K) sex, race, education, years of education, occupation, work class, capital gain, capital loss, work hours per week, relationship, marital status, final weight 1 if sex==0 and race==0 and capital gain + 500 $\cdot$ U > 4000 and years of education + 0.5 $\cdot$ U $\geq$ 14 and work hours per week + U > 40; else 0
@@
 To reflect unmeasured societal influences (e.g., privilege, institutional access), we introduce a latent confounder $U\sim\mathcal{N}(0,1)$, that affects intermediate variables such as years of education, capital gain, and hours of work per week, thereby indirectly influencing income.
+This reflects latent societal advantages that are not captured in observed features.
 The variables influenced by $U$ were selected based on evidence from the fairness literature and sociological theory.
@@
 This models $U$ as an unobserved common cause influencing multiple downstream variables.
+The latent confounder is injected via additive linear terms into structural equations following standard assumptions in structural causal models (SCMs).
+For the binary outcome income, we use a nonlinear threshold function that combines observable and latent contributions.
+This setup reflects bias amplification near decision boundaries and is consistent with prior fairness simulation frameworks.
 Continuous variables are modeled using linear structural equations.
@@
 In our setup, this is instantiated as:
-[…]
-[…]
-[…]
+$\text{Income}=\mathbb{I}\Big[$ $\;\text{capital\_gain}+500\cdot U>4000$
+$\land\;\text{education\_num}+0.5\cdot U\geq 14$
+$\land\;\text{hours\_per\_week}+U>40\Big]$
 This design captures bias amplification, where latent privilege encoded by $U$ influences outcomes near decision boundaries.
@@
 2023).
-Appendix B Appendix: Experimental Details
+A.3 Structured Noise Augmentation
+To test robustness under realistic conditions, we augment the clean synthetic data with structured noise and confounding, emulating sources of measurement error and label corruption common in real-world datasets.
+Additive Noise for Continuous Variables.
+Gaussian noise is added to all continuous variables post-generation to simulate measurement variability.
+For each continuous variable X, noise is sampled as:
+$X_{i}^{\text{noisy}}=X_{i}+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{X}^{2}),\quad\sigma_{X}=\alpha\cdot\mathrm{std}(X)$ (9)
+where $X_{i}$ is the original value for sample $i$, $\epsilon_{i}$ is a random error drawn from a normal distribution with mean 0 and standard deviation $\sigma_{X}$, $\sigma_{X}=\alpha\cdot\mathrm{std}(X)$, where $\alpha\in[0,1]$ is a tunable noise scaling factor set to $\alpha=0.1$.
+This maintains proportionality between noise magnitude and feature variability, consistent with additive noise models used in ICA-LiNGAM and structural equation modeling (SEM) frameworks (Shimizu et al.
+2006; Bollen 1989).
+Symmetric Label Noise in Categorical Variables
+To emulate mislabeling, we introduce symmetric label noise into categorical variables.
+For each categorical variable, a proportion $\gamma$ of values (5%) are randomly replaced with an alternative category, drawn uniformly from the remaining labels.
+For each categorical variable $Y$ with $K$ classes (e.g., ”Male”, ”Female”), uniform symmetric noise is introduced using the following equation.
+Let $\gamma\in[0,1]$ be the label noise probability. Then for each sample $i$:
+[…]
+Where $Y_{i}$ is the original category label, $\tilde{Y}_{i}$ is the observed (possibly noisy) label, $K$ is the number of possible categories for $Y$, and $y^{\prime}\in\mathcal{Y}$ are all possible class values.
+This means with probability $\gamma$, the label is flipped uniformly to one of the K1K-1 incorrect categories.
+And with probability $1-\gamma$, the label remains correct.
+This uniform corruption model known as symmetric label noise, is widely used in label-noise robust learning (Natarajan et al.
+2013; Patrini et al.
+2017).
+Controlled Replication Across Seeds
+To evaluate generalization and fairness robustness, we generate four datasets using different random seeds (42, 101, 222, 333), each with the same underlying graph but distinct noise realizations.
+Noise is added post-graph generation, preserving causal structure while introducing rea)listic variability.
+This mirrors standard practices in fairness benchmarking and robustness testing for CD.
+Appendix B Experimental Details
 In this section, we detail the specific hyperparameters utilized for each set of experiments.
@@
 We perform 200 trials per dataset, optimizing for F1 score via Bayesian search.
-The tuned hyperparameter space includes: query prioritization weights selected from $[(0.3,0.3,0.4),(0.25,0.25,0.5),(0.4,0.3,0.3),(0.3,0.4,0.3)]$, score threshold values between 0.25 and 0.35, temperature values between 0.25 and 0.6, and a maximum iteration range of 40 to 100.
+The tuned hyperparameter space includes: query prioritization weights selected from [(0.3,0.3,0.4),(0.25,0.25,0.5),(0.4,0.3,0.3),(0.3,0.4,0.3)][(0.3,0.3,0.4),(0.25,0.25,0.5),(0.4,0.3,0.3),(0.3,0.4,0.3)], score threshold values between 0.25 and 0.35, temperature values between 0.25 and 0.6, and a maximum iteration range of 40 to 100.
 We report the average across seeds for all performance metrics used in evaluation.
@@
 We focus on paths from sensitive attributes (sex, race, age) to the outcome variable (income) for fairness evaluation.
+An example of a full prompt for this experiment is listed in Listing 2.
+Listing 2 LLM Query Example - Syn-Adult Dataset
+1System: You are an expert in socioeconomic and demographic research.
+2User: You are a helpful assistant to an expert in socioeconomics, with a focus on income prediction and labor statistics.
+3
+4Metadata:
+5Variable A: education: highest level of education achieved (e.g., Bachelors, Masters)
+6Variable B: income: income level (whether the individuals income exceeds \\$50K/year) ...
+7
+8Query:
+9Does education cause income?
+10Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
 Child Network.
@@
 Hyperparameter ranges are: score threshold $[0.01,0.2]$, temperature $[0.05,0.7]$, and max iterations $[10,50]$.
+An example of a full prompt for this experiment is listed in Listing 3.
+Listing 3 LLM Query Example - CHILD Dataset
+1System: You are an expert on childrens diseases.
+2User: You are a helpful assistant to a childrens disease expert.
+3
+4Metadata:
+5Variable A: BirthAsphyxia: lack of oxygen to the blood during the infants birth
+6Variable B: Disease: infant methemoglobinemia
+7
+8Query:
+9Does BirthAsphyxia cause Disease?
+10Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
 Neuropathic Pain Network.
@@
 Parameter ranges include score thresholds $[0.1,0.5]$, temperature $[0.1,0.3]$, and max iterations $[200,500]$.
-Appendix C Appendix: Parameter Impact Analysis on Proposed Method For Child and Neuropathic Networks
+An example of a full prompt for this experiment is listed in Listing 4.
+Listing 4 LLM Query Example - Neuropathic Dataset
+1System: You are an expert on neuropathic pain diagnosis.
+2User: You are a helpful assistant to a neuropathic pain diagnosis expert.
+3
+4Metadata:
+5Variable A: L C6 Radikulopati: left C6 radiculopathy (nerve pain originating at cervical vertebra C6)
+6Variable B: L Handbesvar: left hand discomfort
+7
+8Query:
+9Does L C6 Radikulopati cause L Handbesvar?
+10Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
+Appendix C Parameter Impact Analysis on Proposed Method For Child and Neuropathic Networks
 To evaluate the influence of various parameters on the performance of the proposed method, a detailed analysis of the hyperparameters is conducted using visualizations such as ridge plots, correlation heatmaps, and feature importance analysis.
@@
 These findings underscore the importance of optimizing statistical information and query history weights while treating other parameters as secondary contributors to model performance.
-Appendix D Appendix: Fairness and Path Analyses Results
+Appendix D Fairness and Path Analyses Results
 This table shows the average fairness and path scores for the baseline synthetic graph and the graphs generated by the CD methods.
@@
 PC 0 1 $\pm$ 1 51 $\pm$ 23 0.02 $\pm$ 0.01 0.00 0.221 $\pm$ 0.12 0.221 $\pm$ 0.12 1.326 $\pm$ 1.05
-The graphs generated using the proposed method, PC, GES, DAGMA, Notears, LLM Pairwise, and LLM BFS methods are shown in Figures 10 to 16.
+The graphs generated using the proposed method, PC, GES, DAGMA, Notears, LLM Pairwise, and LLM BFS methods are shown in Figures 9 to 15.
 Figure 9: Causal Graph for Adult-based Synthetic Network Generated by Proposed Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
@@
 Figure 15: Causal Graph for Adult-based Synthetic Network Generated by LLM BFS Method. Red and orange nodes represent target and sensitive attributes respectively, and blue nodes represent others.
-Appendix E Appendix: Complexity Analyses
+Appendix E Complexity Analyses
 The original BFS method (Jiralerspong et al.